### v1

# On the Small Jumps of Lévy Processes and the Multivariate Dickman Distribution

Abstract
We develop a Dickman approximation to the small jumps of multivariate Lévy processes and related stochastic integral processes. Further, we show that the multivariate Dickman distribution is the unique distribution satisfying a certain stability property and that a related stochastic process is self-similar. Along the way we develop applications to non-Gaussian OU-Processes, to the class of generalized multivariate gamma distributions, and to the popular Barndorff-Nielsen and Shephard stochastic volatility model.
1 Introduction
Lévy and related infinitely divisible processes are used in many application areas. When simulating a Lévy process, the difficulty typically lies in the simulation of the small jumps, as the large jumps follow a compound Poisson process, which is often easier to simulate. This has led to the study of approximations of small jumps by simpler processes. In the univariate case, an approximation of small jumps by Brownian motion was developed in [1]. This was extended to the multivariate case in [8], see also the discussion of applications to finance in Chapter 6 of [9]. For important situations where the Brownian motion approximation fails, [10] developed an approximation by a Dickman Lévy process in the univariate case. Recently, a multivariate generalization of the Dickman distribution was introduced in [6] and [15].
The goal of the current paper is three-fold. First, we develop a Dickman approximation to the small jumps of multivariate Lévy processes. Second, we extend these results to stochastic integral processes. Third, we show that the multivariate Dickman distribution is the unique distribution satisfying a certain stability property related to the stability of its jumps under scaling and that a related stochastic process is self-similar. To the best of our knowledge, the second and third of these were previously unknown even in the univariate case. Along the way we develop applications to non-Gaussian processes of Ornstein-Uhlenbeck-type (OU-processes), to the class of generalized multivariate gamma distributions, and to the popular Barndorff-Nielsen and Shephard (BNS) stochastic volatility model introduced in [5].
The rest of this paper is organized as follows. In Section 2 we review infinitely divisible distributions, Lévy processes, and stochastic integral processes and we prove a transfer theorem that allows one to transfer convergence results about sequences of Lévy processes to sequences of stochastic integral processes. This will be needed to prove our main results and may be of independent interest. In Section 3 we recall basic facts about Dickman and multivariate Dickman distributions and introduce the more general class of $\epsilon$-multivariate Dickman distributions. We then show that these are the only distributions that satisfy a certain stability property and that a related stochastic process is self-similar. In Section 4 we give our main results for approximating the small jumps of certain multivariate Lévy processes by Dickman Lévy processes and we extend these results to stochastic integral processes. In Section 5 we give an application to the simulation of generalized multivariate gamma distributions. In Section 6 we give an application to the BNS stochastic volatility model. Proofs are postponed to Section 7.
[…]
$(a,b]C=\left\{x\in\mathbb{R}^{d}:|x|\in(a,b],\frac{x}{|x|}\in C\right\}.$
2 Infinitely Divisible Distributions, Lévy Processes, Stochastic Integrals, and a Transfer Theorem
In which we review basic properties of infinitely divisible distributions and their associated Lévy processes, discuss stochastic integral processes, and prove a transfer theorem that allows one to transfer convergence results about sequences of Lévy processes to sequences of stochastic integral processes. The transfer theorem is needed to prove our main results and may be of independent interest.
The characteristic function of an infinitely divisible distribution $\mu$ on $\mathbb{R}^{d}$ can be written in the form $\hat{\mu}(z)=\exp\{C_{\mu}(z)\}$, where
$C_{\mu}(z)=-\langle z,Az\rangle+i\left\langle b,z\right\rangle+\int_{\mathbb{R}^{d}}\left(e^{i\left\langle z,x\right\rangle}-1-i\left\langle z,x\right\rangle 1_{[|x|\leq 1]}\right)M(\mathrm{d}x),\ \ \ z\in\mathbb{R}^{d},$ (1)
$A$ is a $d\times d$-dimensional covariance matrix called the Gaussian part, $b\in\mathbb{R}^{d}$ is the shift, and $M$ is the Lévy measure, which is a Borel measure on $\mathbb{R}^{d}$ satisfying
$M(\{0\})=0\text{ and }\int_{\mathbb{R}^{d}}(|x|^{2}\wedge 1)M(\mathrm{d}x)<\infty.$
The parameters $A$, $M$, and $b$ uniquely determine this distribution and we write $\mu=\mathrm{ID}(A,M,b)$. We call $C_{\mu}$ the cumulant generating function (cgf) of $\mu$. The normal distribution corresponds to the case where $M=0$ and is denoted $N(b,A)$.
Associated with every infinitely divisible distribution $\mu$ is a Lévy process $\{X_{t}:t\geq 0\}$, where $X_{1}\sim\mu$. Lévy processes are characterized by independent and stationary increments, càdlàg paths, stochastic continuity, and the initial condition $X_{0}=0$ with probability $1$. Standard Brownian motion is the Lévy process $\{W_{t}:t\geq 0\}$ with $W_{1}\sim N(0,I_{d})$, where $I_{d}$ is the $d\times d$-dimensional identity matrix. The jumps of a Lévy process are governed by its Lévy measure. Specifically, if $\{X_{t}:t\geq 0\}$ is a Lévy process with $X_{1}\sim\mathrm{ID}(A,M,b)$, then, for any $B\in\mathfrak{B}(\mathbb{R}^{d})$, $M(B)$ is the expected number of jumps that the process has in the time interval $[0,1]$ that fall inside set $B$.
A Lévy process has finite variation if and only if $A=0$ and $M$ satisfies the additional condition
$\int_{\mathbb{R}^{d}}(|x|\wedge 1)M(\mathrm{d}x)<\infty.$ (2)
Through a slight abuse of terminology, we also say that the associated distribution $\mu$ has finite variation. In this case, the cgf can be written in the form
$C_{\mu}(z)=i\left\langle\gamma,z\right\rangle+\int_{\mathbb{R}^{d}}\left(e^{i\left\langle z,x\right\rangle}-1\right)M(\mathrm{d}x),\ \ \ z\in\mathbb{R}^{d},$ (3)
where $\gamma=b-\int_{|x|\leq 1}xM(\mathrm{d}x)\in\mathbb{R}^{d}$ is the drift, and we write $\mu=\mathrm{ID}_{0}(M,\gamma)$. For more on infinitely divisible distributions and Lévy processes see [27] or [9].
When discussing convergence in distribution of infinitely divisible distributions and Lévy processes, the concept of vague convergence is fundamental. A portmanteau theorem giving several statements that are equivalent to vague convergence can be found in, e.g., [2]. The definition is as follows.
Definition 1.
Let $M_{0},M_{1},M_{2},\dots$ be a sequence of Lévy measures on $\mathbb{R}^{d}$. We say that $M_{n}$ converges vaguely to $M_{0}$ and write $M_{n}\stackrel{{\scriptstyle v}}{{\rightarrow}}M_{0}$ as $n\to\infty$ if $\lim\limits_{n\to\infty}\int_{\mathbb{R}^{d}}f(x)M_{n}(\mathrm{d}x)=\int_{\mathbb{R}^{d}}f(x)M_{0}(\mathrm{d}x)$, for every $f:\mathbb{R}^{d}\mapsto\mathbb{R}$ that is bounded, continuous, and vanishing on a neighborhood of 00.
We now turn to stochastic integral processes. Let ×$\mathbb{R}^{\ell\times d}$ be the space ×$\ell\times d$-dimensional matrices. Let $F(\cdot,\cdot):[0,\infty)^{2}\mapsto\mathbb{R}^{\ell\times d}$ such that $F(\cdot,t)$ is a Borel function for each $t\geq 0$. We write $F_{i,j}(\cdot,\cdot)$ to denote the $i,j$th component of $F$. Assume that, for $t\geq 0$ and each $i,j$,
$\int_{0}^{\infty}|F_{i,j}(s,t)|\vee|F_{i,j}(s,t)|^{2}\mathrm{d}s<\infty.$ (4)
In this case we say that $F$ is an integrable kernel on ×$\mathbb{R}^{\ell\times d}$. Let $Z=\{Z_{t}:t\geq 0\}$ be a $d$-dimensional Lévy process with $Z_{1}\sim\mu=\mathrm{ID}(A,M,b)$ and set
$X_{t}=\int_{0}^{\infty}F(s,t)\mathrm{d}Z_{s},$
where the stochastic integral is defined in the sense of Definition 2.20 in [28]. Many properties of the stochastic integral are given in [28]. We call $X=\{X_{t}:t\geq 0\}$ a stochastic integral process with background driving Lévy process (BDLP) $Z$ and integrable kernel $F$. For each $t\geq 0$, the random variable $X_{t}$ takes values in $\mathbb{R}^{\ell}$. Applying Proposition 2.7 in [28] and dominated convergence, shows that the characteristic function of $X_{t}$ is given by
$\mathrm{E}\left[e^{i\langle z,X_{t}\rangle}\right]=\exp\left\{\int_{0}^{\infty}C_{\mu}(F^{\top}(s,t)z)\mathrm{d}s\right\},\ \ z\in\mathbb{R}^{\ell},$ (5)
where $C_{\mu}$ is the cgf of $Z_{1}$. It follows that $X_{t}\sim\mathrm{ID}(A_{t},M_{t},b_{t})$, where $A_{t}=\int_{0}^{\infty}F(s,t)AF^{\top}(s,t)\mathrm{d}s$,
$M_{t}(B)=\int_{0}^{\infty}\int_{\mathbb{R}^{d}}1_{B\setminus\{0\}}(F(s,t)x)M(\mathrm{d}x)\mathrm{d}s,\ \ B\in\mathfrak{B}(\mathbb{R}^{d}),$
and
$b_{t}=\int_{0}^{\infty}\left(F(s,t)b+\int_{\mathbb{R}^{d}}F(s,t)x\left(1_{[|F(s,t)x|\leq 1]}-1_{[|x|\leq 1]}\right)M(\mathrm{d}x)\right)\mathrm{d}s.$
A common situation is when $\ell=d$ and $F(s,t)=f(s,t)I_{d}$, where $I_{d}$ is the identity matrix in $\mathbb{R}^{d\times d}$ and $f:[0,\infty)^{2}\mapsto\mathbb{R}$ is such that $f(\cdot,t)$ is a Borel function for each $t\geq 0$ satisfying appropriate integrability conditions. In this case we often write
$X_{t}=\int_{0}^{\infty}f(s,t)\mathrm{d}Z_{s}$
instead of $\int_{0}^{\infty}F(s,t)\mathrm{d}Z_{s}$. Perhaps the most famous stochastic integral process is when $f(s,t)=e^{-c(t-s)}1_{[0\leq s\leq t]}$ for some $c>0$. This corresponds to an OU-process, see [9] or [24], and the references therein for details. Another important, although somewhat trivial, example is when $f(s,t)=1_{[0\leq s\leq t]}$. Here the stochastic integral process is the same as the original BDLP.
We now give our transfer theorem. It shows continuity of the stochastic integral process with respect to the BDLP. We have not seen a result of this type in the literature, although related results can be found in [23], [29], Proposition 8.2 in [4], and Proposition 3.7 in [28].
Theorem 1.
For $n=0,1,2,\dots$ let $\{Z_{t}^{(n)}:t\geq 0\}$ be a sequence of Lévy processes on $\mathbb{R}^{d}$ with $Z_{1}^{(n)}\sim\mu^{(n)}=\mathrm{ID}(A^{(n)},M^{(n)},b^{(n)})$ and let $X^{(n)}_{t}=\int_{0}^{\infty}F(s,t)\mathrm{d}Z_{s}^{(n)}$ for $t\geq 0$, where $F$ is an integrable kernel on ×$\mathbb{R}^{\ell\times d}$. Assume that there is a sequence of real numbers $\{c_{n}\}$ with $c_{n}Z_{1}^{(n)}\stackrel{{\scriptstyle d}}{{\rightarrow}}Z_{1}^{(0)}$. If either:
(i) there exists an $h\geq 0$ with $M_{h}:=\limsup_{n\to\infty}|c_{n}|\int_{|c_{n}x|>h}|x|M^{(n)}(\mathrm{d}x)<\infty$, or
(ii) there exist a function $a:[0,\infty)\mapsto[0,\infty)$ with $F(s,t)=0$ for all $s>a(t)$,
then $\{c_{n}X_{t}^{(n)}:t\geq 0\}\stackrel{{\scriptstyle fdd}}{{\rightarrow}}\{X_{t}^{(0)}:t\geq 0\}$.
Note that (ii) always holds when $X_{t}=\int_{0}^{t}F(s,t)\mathrm{d}Z_{s}=\int_{0}^{\infty}1_{[0\leq s\leq t]}F(s,t)\mathrm{d}Z_{s}$. In particular, it always holds for OU-processes.
3 $\epsilon$-Multivariate Dickman Distributions, Stability, and Self-Similarity
In which Dickman and multivariate Dickman distributions are reviewed, the more general class of $\epsilon$-multivariate Dickman distributions is introduced and shown to be characterized by a certain stability property, and a related process is shown to be self-similar.
A random variable $X$ on $\mathbb{R}$ is said to have a generalized Dickman distribution if
$X\stackrel{{\scriptstyle d}}{{=}}U^{1/\theta}(X+1),$
where $\theta>0$ and $U\sim U(0,1)$ is independent of $X$ on the right side. We denote this distribution by $\mathrm{GD}(\theta)$. When $\theta=1$, it is just called the Dickman distribution. Many properties of this distribution are discussed in the surveys [21], [20], and [14]. A multivariate generalization of the Dickman distribution was recently introduced in [6]. It was further studied in [15], where many properties were derived and several approaches for simulation were studied. A random variable $X$ on $\mathbb{R}^{d}$ is said to have a multivariate Dickman (MD) distribution if
$X\stackrel{{\scriptstyle d}}{{=}}U^{1/\theta}(X+\xi),$ (6)
where $\theta>0$ and $X,\xi,U$ are independent on the right side with $U\sim U(0,1)$ and $\xi\sim\sigma_{1}$ for some probability distribution $\sigma_{1}$ on $\mathbb{S}^{d-1}$. Let $\sigma=\theta\sigma_{1}$ and note that $\theta=\sigma(\mathbb{S}^{d-1})$ and $\sigma_{1}=\sigma/\sigma(\mathbb{S}^{d-1})$. Thus, there is no loss of information when working with $\sigma$ instead of $\theta$ and $\sigma_{1}$. We write $\mathrm{MD}(\sigma)$ to denote this distribution. The generalized Dickman distribution $\mathrm{GD}(\theta)$ corresponds to the multivariate Dickman distribution $\mathrm{MD}(\sigma)$ when the dimension $d=1$ and $\sigma=\theta\delta_{1}$. In [15] it is shown that the multivariate Dickman distribution is infinitely divisible. Before giving its Lévy measure, we introduce the more general class of $\epsilon$-multivariate Dickman distributions.
For $\epsilon>0$, an infinitely divisible distribution $\mu$ on $\mathbb{R}^{d}$ is said to be an $\epsilon$-multivariate Dickman distribution if $\mu=\mathrm{ID}_{0}(D^{\epsilon},0)$, where
$D^{\epsilon}(B)=\int_{\mathbb{S}^{d-1}}\int_{0}^{\epsilon}1_{B}(rs)r^{-1}\mathrm{d}r\sigma(\mathrm{d}s),\ \ \ B\in\mathfrak{B}(\mathbb{R}^{d})$ (7)
for some finite Borel measure $\sigma$ on $\mathbb{S}^{d-1}$. We denote this distribution by $\mathrm{MD}^{\epsilon}(\sigma)$ and we refer to $\sigma$ as the spectral measure. The multivariate Dickman distribution corresponds to the case where $\epsilon=1$, i.e. $\mathrm{MD}(\sigma)=\mathrm{MD}^{1}(\sigma)$, see [15]. The characteristic function of $\mu=\mathrm{MD}^{\epsilon}(\sigma)$ is given by
$\hat{\mu}(z)=\exp\left\{\int_{\mathbb{S}^{d-1}}\int_{0}^{\epsilon}\left(e^{ir\left\langle z,s\right\rangle}-1\right)r^{-1}\mathrm{d}r\sigma(\mathrm{d}s)\right\},\ \ \ z\in\mathbb{R}^{d}.$ (8)
Taking partial derivatives of the exponent in (8) shows that the mean vector and the covariance matrix of $X\sim\mathrm{MD}^{\epsilon}(\sigma)$ are given, respectively, by
$\mathrm{E}[X]=\epsilon\int_{\mathbb{S}^{d-1}}s\sigma(\mathrm{d}s)\ \mbox{ and }\ \mathrm{cov}(X)=\frac{\epsilon^{2}}{2}\int_{\mathbb{S}^{d-1}}ss^{\top}\sigma(\mathrm{d}s).$
One reason for the importance of $\epsilon$-multivariate Dickman distributions is that they satisfy a certain stability property, which stems from the following property of the Lévy measure $D^{\epsilon}$ defined in (7). A simple change of variables shows that for any $\epsilon,\delta>0$ and $B\in\mathfrak{B}(\mathbb{R}^{d})$
$D^{1}(B)=D^{\epsilon}(\epsilon B)\ \mbox{ and }\ D^{\epsilon}(\delta B)=D^{\epsilon/\delta}(B).$
It follows that, for any $\epsilon,\gamma>0$, if $X^{\epsilon}\sim\mathrm{MD}^{\epsilon}(\sigma)$, then
$\frac{1}{\gamma}X^{\epsilon}\sim\mathrm{MD}^{\epsilon/\gamma}(\sigma).$ (9)
Remark 1.
Combining (6) with (9) show that $X^{\epsilon}\sim\mathrm{MD}^{\epsilon}(\sigma)$ if and only if
$X^{\epsilon}\stackrel{{\scriptstyle d}}{{=}}U^{1/\theta}(X^{\epsilon}+\epsilon\xi),$
where $\theta=\sigma(\mathbb{S}^{d-1})$ and $X,\xi,U$ are independent on the right side with $U\sim U(0,1)$ and $\xi\sim\sigma/\theta$. In particular, this means that $\epsilon$-multivariate Dickman distributions belong to the class of multivariate Vervaat perpetuities, see [15].
We now show that $\epsilon$-multivariate Dickman distributions are the only ones that satisfy a scaling property like (9).
Theorem 2.
Let $M\neq 0$ be a Borel measure on $\mathbb{R}^{d}$ such that $M(\{0\})=0$ and for each $h\in(0,\infty)$
$\int_{|x|\leq h}|x|^{2}M(\mathrm{d}x)<\infty.$
Let $M^{\delta}(\mathrm{d}x)=1_{[|x|\leq\delta]}M(\mathrm{d}x)$ for any $\delta>0$. This is a Lévy measure for every $\delta>0$ and we write $X^{\delta}$ to denote a random variable with distribution $\mathrm{ID}(0,M^{\delta},0)$. Assume that there exists a function $\psi:(0,\infty)\mapsto(0,\infty)$ such that for any $\epsilon,\gamma>0$ there exists a non-random $b_{*}(\epsilon,\gamma)\in\mathbb{R}^{d}$ with
$\frac{1}{\psi(\gamma)}X^{\epsilon}\stackrel{{\scriptstyle d}}{{=}}X^{\epsilon/\gamma}+b_{*}(\epsilon,\gamma).$ (10)
Then $\psi(\gamma)=\gamma$ and there exists a spectral measure $\sigma\neq 0$ and a $b_{\epsilon}\in\mathbb{R}^{d}$ such that $X^{\epsilon}+b_{\epsilon}\sim\mathrm{MD}^{\epsilon}(\sigma)$.
Note that a measure $M$ satisfies the assumption of Theorem 2 if and only if $M^{\delta}(\mathrm{d}x)=1_{[|x|\leq\delta]}M(\mathrm{d}x)$ is a Lévy measure for each $\delta>0$. We now show the self-similarity of a process related to $\epsilon$-multivariate Dickman distributions. Let $\sigma$ be a finite Borel measure on $\mathbb{S}^{d-1}$, let $\theta=\sigma(\mathbb{S}^{d-1})$, let $E_{1},E_{2},\dots\stackrel{{\scriptstyle\mathrm{iid}}}{{\sim}}\mathrm{Exp}(\theta)$, let $\Gamma_{i}=\sum_{k=1}^{i}E_{i}$, and let $\xi_{1},\xi_{2},\dots\stackrel{{\scriptstyle\mathrm{iid}}}{{\sim}}\sigma_{1}:=\sigma/\theta$ be independent of the $E_{i}$’s. In [15] it is shown that
$\sum_{i=1}^{\infty}e^{-\Gamma_{i}}\xi_{i}\sim\mathrm{MD}(\sigma).$ (11)
This can be used to simulate from a multivariate Dickman distribution. However, to do so we must truncate the infinite series. It has been argued, see e.g. [9], that the proper way to truncate is not at a deterministic point. Instead, we should choose some $\tau>0$ and truncate at the random time $N(\tau)$, where $N(\tau)=\max\{i:\Gamma_{i}\leq\tau\}$. This way we have control on the magnitudes of the jumps that we are removing. The remainder in this approximation can be written as
$R_{\tau}=\sum_{i=N(\tau)+1}^{\infty}e^{-\Gamma_{i}}\xi_{i}=\sum_{i=1}^{\infty}1_{[\Gamma_{i}>\tau]}e^{-\Gamma_{i}}\xi_{i}.$
We select $\tau=\tau(t)=\log(1/t)$ for $t\in[0,1]$ and set
$X_{t}=R_{\tau(t)}=\sum_{i=1}^{\infty}1_{[\Gamma_{i}>\log(1/t)]}e^{-\Gamma_{i}}\xi_{i}$ (12)
to be the remainder process. Here we have $X_{0}=0$ a.s. The reason for our choice of $\tau$ is that $\Gamma_{i}>\log(1/t)$ is equivalent to $e^{-\Gamma_{i}}<t$. We now show that $X=\{X_{t}:t\in[0,1]\}$ is a self-similar process.
Theorem 3.
Let $X_{t}$ be as in (12). We have $X_{t}\sim\mathrm{MD}^{t}(\sigma)$ and for any $a\in[0,1]$, we have
$\{X_{at}:t\in[0,1]\}\stackrel{{\scriptstyle d}}{{=}}\{aX_{t}:t\in[0,1]\}.$
4 Approximating Small Jumps of Lévy processes
In which a limit theorem showing that the small jumps of a large class of Lévy processes can be well approximated by a Dickman Lévy process is proved and a related limit theorem for stochastic integral processes is given.
Here and throughout, when applied to Lévy processes $\stackrel{{\scriptstyle d}}{{\rightarrow}}$ refers to weak convergence on the space $D([0,\infty),\mathbb{R}^{d})$, which is the space of càdlàg functions from $[0,\infty)$ into $\mathbb{R}^{d}$ equipped with the Skorokhod topology. For any $\gamma\in\mathbb{R}^{d}$, we write $\gamma^{*}=\{t\gamma:t\geq 0\}$ to denote the element of $D([0,\infty),\mathbb{R}^{d})$ that maps $t$ to $t\gamma$. In particular, $0^{*}$ denotes the function that is identically zero.
Let $X=\{X_{t}:t\geq 0\}$ be a Lévy process with $X_{1}\sim\mathrm{ID}_{0}(\nu,0)$. Note that here, for simplicity, we set the drift to zero. Fix $\epsilon>0$ and consider the truncated Lévy process $X^{\epsilon}=\{X_{t}^{\epsilon}:t\geq 0\}$ obtained by removing the jumps of the process $X$, whose magnitudes exceed $\epsilon$. In this case, $X^{\epsilon}_{1}\sim\mathrm{ID}_{0}(\nu^{\epsilon},0)$, where $\nu^{\epsilon}(B)=\int_{|x|\leq\epsilon}1_{B}(x)\nu(\mathrm{d}x)$, $B\in\mathfrak{B}(\mathbb{R}^{d})$. Next, consider the scaled truncated process $\epsilon^{-1}X^{\epsilon}=\{\epsilon^{-1}X_{t}^{\epsilon}:t\geq 0\}$ and note that all of its jumps are bounded by $1$. It is easily checked that $\epsilon^{-1}X_{1}^{\epsilon}\sim\mathrm{ID}_{0}(M^{\epsilon},0)$, where
$M^{\epsilon}(B)=\int_{|x|\leq\epsilon}1_{B}\left(\frac{x}{\epsilon}\right)\nu(\mathrm{d}x)=\int_{\mathbb{R}^{d}}1_{B}\left(\frac{x}{\epsilon}\right)\nu^{\epsilon}(\mathrm{d}x)=\nu^{\epsilon}(\epsilon B),\ \ \ B\in\mathfrak{B}(\mathbb{R}^{d}).$
Now, consider the multivariate Dickman Lévy process $Y^{1}=\{Y_{t}^{1}:t\geq 0\}$ with $Y_{1}^{1}\sim\mathrm{MD}(\sigma)$. We now give conditions for the scaled truncated process $\epsilon^{-1}X^{\epsilon}$ to converge to $Y^{1}$ in distribution. For the convergence to hold, we need $M^{\epsilon}\stackrel{{\scriptstyle v}}{{\rightarrow}}D^{1}$ as $\epsilon\downarrow 0$, where $M^{\epsilon}$ is the Lévy measure of $\epsilon^{-1}X^{\epsilon}_{1}$ and $D^{1}$ is the Lévy measure of $Y^{1}_{1}$. We now give several statement that are equivalent to this. In the univariate case, a version of this result is given in Proposition 2.1 of [10]. As usual, for a set $C\in\mathfrak{B}(\mathbb{S}^{d-1})$, we write $\partial C$ to denote its boundary.
Proposition 1.
Let $M^{\epsilon}$, $D^{1}$, and $\sigma$ be as described above. The following statements are equivalent:
  • 1.
    $M^{\epsilon}\stackrel{{\scriptstyle v}}{{\rightarrow}}D^{1}$ as $\epsilon\downarrow 0$.
  • 2.
    For all $10<h<$ 1 and all $C\in\mathfrak{B}(\mathbb{S}^{d-1})$ with $\sigma(\partial C)=0$, $\nu((\epsilon h,\epsilon]C)\rightarrow\sigma(C)\log\frac{1}{h}$ as $\epsilon\downarrow 0$.
  • 3.
    For all $p>0$ and all $C\in\mathfrak{B}(\mathbb{S}^{d-1})$ with $\sigma(\partial C)=0$, $\frac{1}{\epsilon^{p}}\int_{(0,\epsilon]C}|x|^{p}\nu(\mathrm{d}x)\rightarrow\frac{\sigma(C)}{p}$ as $\epsilon\downarrow 0$.
  • 4.
    For some $p>0$ and all $C\in\mathfrak{B}(\mathbb{S}^{d-1})$ with $\sigma(\partial C)=0$, $\frac{1}{\epsilon^{p}}\int_{(0,\epsilon]C}|x|^{p}\nu(\mathrm{d}x)\rightarrow\frac{\sigma(C)}{p}$ as $\epsilon\downarrow 0$.
Note that we allow $\sigma=0$ in the above. Part of the result is the fact that, so long as any of the equivalent conditions in Proposition 1 hold, for every $p>0$ we have
$\int_{(0,\epsilon]C}|x|^{p}\nu(\mathrm{d}x)<\infty.$
We now give our main result for approximating the small jumps of Lévy processes by a Dickman Lévy process. This extends the univariate result in [10] to the multivariate case.
Theorem 4.
We have $\epsilon^{-1}{X^{\epsilon}}\stackrel{{\scriptstyle d}}{{\rightarrow}}Y^{1}$ as $\epsilon\downarrow 0$ if and only if any of the equivalent conditions in Proposition 1 hold. Further, when $\sigma(\mathbb{S}^{d-1})=0$, this is equivalent to $\epsilon^{-1}X^{\epsilon}\stackrel{{\scriptstyle d}}{{\rightarrow}}0^{*}$ as $\epsilon\downarrow 0$, and when $\sigma(\mathbb{S}^{d-1})>0$, it is equivalent to
$\left(\frac{\sigma(\mathbb{S}^{d-1})}{p\int_{|x|\leq\epsilon}|x|^{p}\nu(\mathrm{d}x)}\right)^{1/p}X^{\epsilon}\stackrel{{\scriptstyle d}}{{\rightarrow}}Y^{1}\mbox{ as }\epsilon\downarrow 0$
for any $p>0$.
We now give conditions that are easily checked in an important special case. Let $L^{1}(\mathbb{S}^{d-1},\sigma)$ be the space of Borel functions $g:\mathbb{S}^{d-1}\mapsto\mathbb{R}$ with $\int_{\mathbb{S}^{d-1}}|g(s)|\sigma(\mathrm{d}s)<\infty$. For any nonnegative $g\in L^{1}(\mathbb{S}^{d-1},\sigma)$ define the finite measure $\sigma_{g}$ by $\sigma_{g}(B)=\int_{B}g(s)\sigma(\mathrm{d}s)$ for $B\in\mathfrak{B}(\mathbb{S}^{d-1})$.
Corollary 1.
Assume that $\nu$ is of the form
$\nu(B)=\int_{\mathbb{S}^{d-1}}\int_{0}^{\infty}1_{B}(rs)\rho(r,s)\mathrm{d}r\sigma(\mathrm{d}s),\ \ \ B\in\mathfrak{B}(\mathbb{S}^{d-1}),$
where $\sigma$ is a finite Borel measure on $\mathbb{S}^{d-1}$ and $\rho:[0,\infty)\times\mathbb{S}^{d-1}\mapsto[0,\infty)$ is a Borel function. If there is a nonnegative $g\in L^{1}(\mathbb{S}^{d-1},\sigma)$ with $r\rho(r,\cdot)\to g(\cdot)$ in $L^{1}(\mathbb{S}^{d-1},\sigma)$ as $r\downarrow 0$, i.e. $\int_{\mathbb{S}^{d-1}}|r\rho(r,s)-g(s)|\sigma(\mathrm{d}s)\to 0$ as $r\downarrow 0$, then the equivalent conditions in Proposition 1 hold with $\sigma_{g}$ in place of $\sigma$ and $\frac{X^{\epsilon}}{\epsilon}\stackrel{{\scriptstyle d}}{{\rightarrow}}Y^{1}$ as $\epsilon\downarrow 0$, where $Y^{1}_{1}\sim\mathrm{MD}(\sigma_{g})$.
We now give two ways of checking that convergence in $L^{1}(\mathbb{S}^{d-1},\sigma)$ holds.
Remark 2.
Assume that, for each $s\in\mathbb{S}^{d-1}$, $r\rho(r,s)\to g(s)$ as $r\downarrow 0$.
1. If $g$ is bounded and $r\rho(r,s)$ is bounded for small enough $r$, then we have convergence in $L^{1}(\mathbb{S}^{d-1},\sigma)$. This follows by dominated convergence and the fact that $\sigma$ is a finite measure.
2. If $r\rho(r,s)\to g(s)$ uniformly in $s$, then we have convergence in $L^{1}(\mathbb{S}^{d-1},\sigma)$. This is trivial when $\sigma=0$. To see that it holds when $\sigma\neq 0$ note that, in this case, for every $\epsilon>0$ there exists a $\delta>0$ such that for every $s\in\mathbb{S}^{d-1}$ and every $r\in(0,\delta)$ we have $|r\rho(r,s)-g(s)|<\frac{\epsilon}{\sigma(\mathbb{S}^{d-1})}$. Thus, for such $r$, $\int_{\mathbb{S}^{d-1}}|r\rho(r,s)-g(s)|\sigma(\mathrm{d}s)\leq\int_{\mathbb{S}^{d-1}}\frac{\epsilon}{\sigma(\mathbb{S}^{d-1})}\sigma(\mathrm{d}s)=\epsilon$.
Example 1.
In [25] the class of so-called generalized tempered stable distributions was introduced. This class consists of infinitely divisible distributions with no Gaussian part and a Lévy measure of the form
$\nu(B)=\int_{\mathbb{S}^{d-1}}\int_{0}^{\infty}1_{B}(rs)q(r,s)r^{-1-\alpha}\mathrm{d}r\sigma(\mathrm{d}s),\ \ \ B\in\mathfrak{B}(\mathbb{S}^{d-1}),$
where $\alpha\in(0,2)$, $\sigma$ is a finite measure on $\mathbb{S}^{d-1}$, and $q:[0,\infty)\times\mathbb{S}^{d-1}\mapsto(0,\infty)$ is a Borel function such that $q(r,\cdot)\to q_{*}(\cdot)$ in $L^{1}(\mathbb{S}^{d-1},\sigma)$ as $r\downarrow 0$ for some nonnegative $q_{*}\in L^{1}(\mathbb{S}^{d-1},\sigma)$. In the limiting case when $\alpha=0$, these immediately satisfy the assumptions of Corollary 1. This class contains all gamma distributions and many of their extensions, see Section 5 below for a detailed discussion.
Theorem 4 tells us that, under appropriate conditions, we can approximately simulate from a Lévy process by approximating the small jumps by a Dickman Lévy process and the large jumps by a compound Poisson process. To see this let $X=\{X_{t}:t\geq 0\}$ be a Lévy process with $X_{1}\sim\mathrm{ID}_{0}(\nu,\gamma)$. Note that we now allow for a nonzero drift. Set $\epsilon>0$ and let $\nu^{\epsilon}(B)=\int_{B}1_{|x|\leq\epsilon}(x)\nu(\mathrm{d}x)$ and $\tilde{\nu}^{\epsilon}(B)=\int_{B}1_{|x|>\epsilon}(x)\nu(\mathrm{d}x)$ for every $B\in\mathfrak{B}(\mathbb{R}^{d})$. It follows that
$X\stackrel{{\scriptstyle d}}{{=}}X^{\epsilon}+\tilde{X}^{\epsilon}+\gamma^{*},$ (13)
where $X^{\epsilon}=\{X^{\epsilon}_{t}:t\geq 0\}$ is a Lévy process with $X_{1}\sim\mathrm{ID}_{0}(\nu^{\epsilon},0)$, $\tilde{X}^{\epsilon}=\{\tilde{X}^{\epsilon}_{t}:t\geq 0\}$ is a Lévy process with $\tilde{X}_{1}\sim\mathrm{ID}_{0}(\tilde{\nu}^{\epsilon},0)$, $X^{\epsilon}$ and $\tilde{X}^{\epsilon}$ are independent, and $\gamma^{*}=\{t\gamma:t\geq 0\}$. Here, we separated $X$ into: $X^{\epsilon}$, the process of small jumps, $\tilde{X}^{\epsilon}$, the process of large jumps, and $\gamma^{*}$, the deterministic drift process. Under appropriate conditions, when $\epsilon>0$ is small, we can approximate $X^{\epsilon}$ by $\epsilon$ times a Dickman Lévy process $Y^{1}$, which gives
$X\stackrel{{\scriptstyle d}}{{\approx}}\epsilon Y^{1}+\tilde{X}^{\epsilon}+\gamma^{*},$ (14)
where $Y^{1}$ is independent of $\tilde{X}^{\epsilon}$. Following ideas in [8] about approximations of small jumps by Brownian motion, we now verify that the error in this approximation approaches 00 as $\epsilon\downarrow 0$.
Corollary 2.
Let $\{X_{t}:t\geq 0\}$ be a Lévy process on $\mathbb{R}^{d}$ with $X_{1}\sim\mathrm{ID}_{0}(\nu,\gamma)$ and let $Y^{1}$, $\tilde{X}^{\epsilon}$, and $\gamma^{*}$ be as above. If any of the equivalent conditions in Proposition 1 hold, then for every $\epsilon>0$, there exists a càdlàg process $Z^{\epsilon}=\{Z_{t}^{\epsilon}:t\geq 0\}$ such that
$X\stackrel{{\scriptstyle d}}{{=}}\epsilon Y^{1}+\tilde{X}^{\epsilon}+\gamma^{*}+Z^{\epsilon},$
and, for each $T>0$,
$\sup\limits_{t\in[0,T]}|\epsilon^{-1}Z_{t}^{\epsilon}|\stackrel{{\scriptstyle p}}{{\rightarrow}}0\ \text{as}\ \epsilon\downarrow 0.$ (15)
This suggest that, for small $\epsilon$, we can approximate $Z^{\epsilon}$ by 00 and we can approximately simulate from $X$ by independently simulating $Y^{1}$ and $\tilde{X}^{\epsilon}$ and then applying (14). Several methods for simulating $Y^{1}$ are discussed in [15]. For instance, we can simulate up to a finite time horizon $T>0$ as follows. Let $E_{1},E_{2},\dots\stackrel{{\scriptstyle\mathrm{iid}}}{{\sim}}\mathrm{Exp}(\theta)$, let $U_{1},U_{2},\dots\stackrel{{\scriptstyle\mathrm{iid}}}{{\sim}}U(0,1)$, and let $\xi_{1},\xi_{2},\dots\stackrel{{\scriptstyle\mathrm{iid}}}{{\sim}}\sigma/\theta$, where $\theta=\sigma(\mathbb{S}^{d-1})$, be independent sequences of random variables. If $\Gamma_{i}=\sum_{k=1}^{i}E_{i}$ for $i=1,2,\dots$, then
$\left\{Y^{1}_{t}:0\leq t\leq T\right\}\stackrel{{\scriptstyle d}}{{=}}\left\{\sum_{i=1}^{\infty}e^{-\Gamma_{i}/T}\xi_{i}1_{[0,t/T]}(U_{i}):0\leq t\leq T\right\}.$ (16)
In practice, of course, the infinite sum must be truncated at some finite (possibly random) value. We now turn to the simulation of $\tilde{X}^{\epsilon}$. Toward this end, note that $\tilde{\nu}^{\epsilon}$ is a finite measure, set $\lambda^{\epsilon}=\tilde{\nu}^{\epsilon}(\mathbb{R}^{d})$, and let $\tilde{\nu}^{\epsilon}_{p}=\tilde{\nu}^{\epsilon}/\lambda^{\epsilon}$. It follows that $\tilde{\nu}^{\epsilon}_{p}$ is a probability measure. Let $W_{1},W_{2},\dots\stackrel{{\scriptstyle\mathrm{iid}}}{{\sim}}\tilde{\nu}^{\epsilon}_{p}$ and, independent of this, let $\{N(t):t\geq 0\}$ be a Poisson process with rate $\lambda^{\epsilon}$. It is readily checked that $\tilde{X}^{\epsilon}$ is a compound Poisson process and that
$\tilde{X}^{\epsilon}=\left\{\tilde{X}_{t}^{\epsilon}:t\geq 0\right\}\stackrel{{\scriptstyle d}}{{=}}\left\{\sum_{i=1}^{N(t)}W_{i}:t\geq 0\right\}.$ (17)
When simulating until a finite time horizon $T>0$, it is often more convenient to use the following representation. Let $W_{1},W_{2},\dots\stackrel{{\scriptstyle\mathrm{iid}}}{{\sim}}\tilde{\nu}^{\epsilon}_{p}$ and $U_{1},U_{2},\dots\stackrel{{\scriptstyle\mathrm{iid}}}{{\sim}}U(0,1)$ be independent sequences, and, independent of these, let $N\sim\mathrm{Pois}(T\lambda_{\epsilon})$. We have
$\left\{\tilde{X}_{t}^{\epsilon}:0\leq t\leq T\right\}\stackrel{{\scriptstyle d}}{{=}}\left\{\sum_{i=1}^{N}W_{i}1_{[0,t/T]}(U_{i}):0\leq t\leq T\right\},$ (18)
see, e.g., Section 6.1 in [9] for details. Thus, the problem reduces to simulation from $\tilde{\nu}^{\epsilon}_{p}$. In the next section we discuss this in an important situation. Now, we give a version of Theorem 4 for stochastic integral processes. The proof is based on combining Theorems 1 and 4.
Theorem 5.
Let $X^{\epsilon}$ and $Y^{1}$ be as in Theorem 4 and let $F$ be any integrable kernel on ×$\mathbb{R}^{\ell\times d}$. If any of the equivalent condition in Proposition 1 hold, then
$\left\{\frac{1}{\epsilon}\int_{0}^{\infty}F(s,t)\mathrm{d}X^{\epsilon}_{s}:t\geq 0\right\}\stackrel{{\scriptstyle fdd}}{{\rightarrow}}\left\{\int_{0}^{\infty}F(s,t)\mathrm{d}Y_{s}:t\geq 0\right\}\ \ \mbox{as }\ \epsilon\downarrow 0.$
In particular, this holds for OU-processes.
5 Simulation from General Multivariate Gamma Distributions
In which Theorem 4 is applied to develop an approximate simulation method for a large class of distributions and a small simulation study is given to verify the performance of this approximation.
In this section we focus on distributions of the form $\mu=\mathrm{ID}_{0}(\nu,\gamma)$, where
$\nu(B)=\int_{\mathbb{S}^{d-1}}\int_{0}^{\infty}1_{B}(rs)q(r^{p},s)r^{-1}\mathrm{d}r\sigma(\mathrm{d}s),\ \ \ B\in\mathfrak{B}(\mathbb{R}^{d}),$ (19)
$p>0$, $\sigma$ is a finite Borel measure on $\mathbb{S}^{d-1}$, and $q:(0,\infty)\times\mathbb{S}^{d-1}\mapsto[0,\infty)$ is a Borel function such that for each $s\in\mathbb{S}^{d-1}$ we have $\lim_{r\downarrow 0}q(r,s)=1$ and $q(\cdot,s)$ is completely monotone. In [11] such distributions are called proper $p$-tempered 00-stable distributions, see also [12] for many properties. The complete monotonicity of $q$ along with our other assumptions implies that $q(r^{p},s)=\int_{(0,\infty)}e^{-r^{p}s}Q_{s}(\mathrm{d}v)$ for some measurable family $\{Q_{s}\}_{s\in\mathbb{S}^{d-1}}$ of probability measures on $(0,\infty)$ satisfying
$\int_{\mathbb{S}^{d-1}}\int_{0}^{1}\log(1/v)Q_{s}(\mathrm{d}v)\sigma(\mathrm{d}s)<\infty,$
see Remark 1 and Corollary 1 in [11]. If we take $p=1$ and $Q_{s}=\delta_{b(s)}$ for some Borel function $b:\mathbb{S}^{d-1}\mapsto(0,\infty)$, these distributions reduce to the class of multivariate gamma distributions introduced in [22]. Motivated by this, we refer to our class as the class of generalized multivariate gamma distributions (GMGD).
Remark 3.
We can consider a slight generalization of GMGD. Specifically, we can allow for a Gaussian part and remove the requirement that $\lim_{r\downarrow 0}q(r,s)=1$ holds for each $s\in\mathbb{S}^{d-1}$. The resulting distributions are denoted $J_{0,p}$ in [18] and are called extended $p$-tempered 00-stable distributions in [12]. When $p=2$ they are called class $M$, see [17] and the references therein. When $p=1$ we get the Thorin class, which is the smallest class of distributions on $\mathbb{R}^{d}$ that is closed under convolution and weak convergence and contains the distributions of all so-called elementary gamma random variables on $\mathbb{R}^{d}$, see [3]. A related characterization for every $p>0$ is given in Theorem 4.18 of [12].
Note that $q(r^{p},s)\leq 1$, thus Remark 2 guarantees that the assumptions of Theorem 4 and Corollary 2 are satisfied and that these distributions fall into the more general class discussed in Example 1. This means that we can approximately simulate from the GMGD Lévy process $X=\{X_{t}:t\geq 0\}$ with $X_{1}\sim\mathrm{ID}_{0}(\nu,\gamma)$ and $\nu$ as in (19) by using the approximation in (14). Here, to simulate $\tilde{X}^{\epsilon}=\{\tilde{X}^{\epsilon}_{t}:t\geq 0\},$ the process of large jumps, we can either use (17) or (18). Either way, we need a way to simulate from $\tilde{\nu}^{\epsilon}_{p}$. We now develop an approach to do this.
Fix $\epsilon>0$, let $\tilde{\nu}^{\epsilon}$ be the finite Borel measure defined by
$\tilde{\nu}^{\epsilon}(B)=\int_{\mathbb{S}^{d-1}}\int_{\epsilon}^{\infty}1_{B}(rs)q(r^{p},s)r^{-1}\mathrm{d}r\sigma(\mathrm{d}s),\ \ \ B\in\mathfrak{B}(\mathbb{R}^{d}),$
let $\lambda^{\epsilon}=\tilde{\nu}^{\epsilon}(\mathbb{R}^{d})$, and let $\tilde{\nu}^{\epsilon}_{p}=\tilde{\nu}^{\epsilon}/\lambda^{\epsilon}$ be a probability measure. Our goal is to develop an approach to simulate from $\tilde{\nu}^{\epsilon}_{p}$. We begin by defining several quantities and distributions. First, for $u>0$ define
$\ell(u)=\int_{u}^{\infty}r^{-1}e^{-r^{p}}\mathrm{d}r=\frac{1}{p}\int_{u^{p}}^{\infty}y^{-1}e^{-y}\mathrm{d}y=\frac{1}{p}\Gamma(0,u^{p}),$
where $\Gamma(\cdot,\cdot)$ is the upper incomplete gamma function. Next, for $s\in\mathbb{S}^{d-1}$, define
$k_{\epsilon}(s)$ $=$ $\int_{\epsilon}^{\infty}q(r^{p},s)r^{-1}\mathrm{d}r$
$=$ $\int_{\epsilon}^{\infty}\int_{0}^{\infty}r^{-1}e^{-r^{p}v}Q_{s}(\mathrm{d}v)\mathrm{d}r$
$=$ $\frac{1}{p}\int_{0}^{\infty}\int_{\epsilon v^{1/p}}^{\infty}r^{-1}e^{-r^{p}}\mathrm{d}rQ_{s}(\mathrm{d}v)=\int_{0}^{\infty}\ell(\epsilon v^{1/p})Q_{s}(\mathrm{d}v).$
Now, define a probability measure on $\mathbb{S}^{d-1}$ by
$\sigma_{p}(\mathrm{d}s)=\frac{k_{\epsilon}(s)}{\lambda^{\epsilon}}\sigma(\mathrm{d}s)$
and a family of probability measures on $(0,\infty)$ by
$G_{V}(\mathrm{d}v;s)=\frac{\ell(\epsilon v^{1/p})}{k_{\epsilon}(s)}Q_{s}(\mathrm{d}v),$
where $s\in\mathbb{S}^{d-1}$ is a parameter. Finally, define a family of probability measures on $(0,\infty)$ by $G_{R}(\mathrm{d}r;a)=g_{R}(r;a)\mathrm{d}r$, where $a>0$ is a parameter and
$g_{R}(r;a)=\frac{1}{\ell(a)}r^{-1}e^{-r^{p}}1_{[r\geq a]}$ (20)
is the probability density function (pdf).
Proposition 2.
Let $S\sim\sigma_{p}$. Given $S$, let $V\sim G_{V}(\cdot;S)$ and, given $S$ and $V$, let $R\sim G_{R}(\cdot;V^{1/p}\epsilon)$. If $W=RV^{-1/p}S$, then $W\sim\tilde{\nu}^{\epsilon}_{p}$.
There is no general approach for simulating from $\sigma_{p}$ and $G_{V}$ as they depend on the measures $\sigma$ and $Q_{s}$. However, see [13], [30], and the references therein for discussions of simulation from distributions on $\mathbb{S}^{d-1}$. We now develop a rejection sampling method to simulate from $G_{R}$. Toward this end we introduce two distributions that will serve as our proposal distributions. Let
$h_{1}(x;a,p)=px^{p-1}e^{a^{p}-x^{p}}1_{[x\geq a]}$
be a pdf, where $a,p>0$ are parameters. Next, let
$h_{2}(x;a,p,\beta)=\beta\frac{x^{-1}}{\log(1/a)}1_{[a\leq x<1]}+(1-\beta)px^{p-1}e^{1-x^{p}}1_{[x\geq 1]}$
be a pdf, where $a,\beta\in(0,1)$, and $p>0$ are parameters. We can simulate from these distributions as follows.
Lemma 1.
Fix $a>0$, $p>0$, $\beta\in(0,1)$, and let $U\sim U(0,1)$. We have
$\left(a^{p}-\log(U)\right)^{\frac{1}{p}}\sim h_{1}(\cdot;a,p).$
Further, if $a\in(0,1)$, then
$a^{(1-U/\beta)}1_{[U\leq\beta]}+\left(1-\log(1-U)+\log(1-\beta)\right)^{\frac{1}{p}}1_{[U>\beta]}\sim h_{2}(\cdot;a,p,\beta).$
If $a\geq 1$, we can check that
$g_{R}(x)\leq C_{1}h_{1}(x;a,p),\ \mbox{ where }\ C_{1}=\frac{1}{e^{a^{p}}p\ell(a)}.$
Similarly, if $a\in(0,1)$, we can check that
$g_{R}(x)\leq C_{2}h_{2}(x;a,p,\beta),\ \mbox{ where }\ C_{2}=\frac{1}{\ell(a)}\max\left\{\frac{1}{ep(1-\beta)},\frac{\log(1/a)}{\beta}\right\}.$
From here we can derive the following rejection sampling algorithms. Let
$\phi_{1}(x)=x^{-1}1_{[x\geq a]}$
and let
$\phi_{2}(x)=\frac{1}{\max\left\{\frac{1}{ep(1-\beta)},\frac{\log(1/a)}{\beta}\right\}\left(\beta\frac{e^{x^{p}}}{\log(1/a)}1_{[a\leq x<1]}+(1-\beta)epx^{p}1_{[x\geq 1]}\right)}.$
Algorithm 1: Simulation from $G_{R}$ when $p>0$ and $a\geq 1$.
Step 1. Simulate $U_{1},U_{2}\stackrel{{\scriptstyle\mathrm{iid}}}{{\sim}}U(0,1)$ and set $X=\left(a^{p}-\log(U_{1})\right)$.
Step 2. If $U_{2}\leq\phi_{1}(X)$ return $X^{1/p}$, otherwise go back to Step 1.
In this case, the probability of rejection on a given iteration is $1/C_{1}=e^{a^{p}}p\ell(a)$.
Algorithm 2: Simulation from $G_{R}$ when $p>0$ and $a\in(0,1)$; $\beta\in(0,1)$ is a tuning parameter.
Step 1. Simulate $U_{1},U_{2}\stackrel{{\scriptstyle\mathrm{iid}}}{{\sim}}U(0,1)$ and set
$X=a^{1-U_{1}/\beta}1_{[U_{1}\leq\beta]}+\left(1-\log(1-U_{1})+\log(1-\beta)\right)^{\frac{1}{p}}1_{[U_{1}>\beta]}.$
Step 2. If $U_{2}\leq\phi_{2}(X)$ return $X$, otherwise go back to Step 1.
In this case, the probability of rejection on a given iteration is $1/C_{2}$. We will generally take $a=\epsilon V^{1/p}$ with $\epsilon$ small. Thus, we are most interested in the case when $a\to 0$. By l’Hôpital’s rule, we have
$\lim_{a\to 0}\frac{1}{C_{2}}=\beta\lim_{a\to 0}\frac{\ell(a)}{\log(1/a)}=\beta\lim_{a\to 0}\frac{\int_{a}^{\infty}r^{-1}e^{-r^{p}}\mathrm{d}r}{-\log(a)}=\beta\lim_{a\to 0}\frac{-a^{-1}e^{-a^{p}}}{-a^{-1}}=\beta.$
Thus, for small $a$ we can select a large $\beta$ to get a good performance. In general, one can select whichever value of $\beta$ maximizes the acceptance probability in a given situation. For simplicity, throughout this paper we take $\beta=1/2$, which leads to a reasonable performance in the situations considered. We now summarize our algorithm to simulate $\tilde{X}^{\epsilon}$, the compound Poisson process of large jumps. It combines (18) with Algorithms 1 and 2.
Algorithm 3: For $\epsilon>0$, simulate $\tilde{X}^{\epsilon}$, the compound Poisson process of large jumps, up to time $T>0$. Fix the tuning parameter $\beta\in(0,1)$.
  1. I.
    Simulate $N\sim\mathrm{Pois}(T\lambda^{\epsilon})$.
  2. II.
    Simulate $U_{1},U_{2},\dots,U_{N}\stackrel{{\scriptstyle\mathrm{iid}}}{{\sim}}U(0,1)$.
  3. III.
    For $i=1,2,\dots,N$:
    1. 1.
      Simulate $S_{i}\sim\sigma_{p}$.
    2. 2.
      Given $S_{i}$, simulate $V_{i}\sim G_{V}(\cdot;S)$.
    3. 3.
      Given $S_{i}$ and $V_{i}$, simulate $R_{i}\sim G_{R}(\cdot;V^{1/p}\epsilon)$ as follows:
      1. (a)
        If $\epsilon V_{i}^{1/p}\geq 1$:
        1. (a.1)
          Generate $U^{\prime}_{1},U^{\prime}_{2}\stackrel{{\scriptstyle\mathrm{iid}}}{{\sim}}U(0,1)$ and set $X_{i}=\left(\epsilon^{p}V_{i}-\log U^{\prime}_{1}\right)$.
        2. (a.2)
          If $U^{\prime}_{2}\leq X_{i}^{-1}$ set $R_{i}=X_{i}^{1/p}$, otherwise go back to (a.1).
      2. (b)
        If $\epsilon V_{i}^{\frac{1}{p}}<1$:
        1. (b.1)
          Generate $U^{\prime}_{1},U^{\prime}_{2}\stackrel{{\scriptstyle\mathrm{iid}}}{{\sim}}U(0,1)$.
        2. (b.2)
          If $U^{\prime}_{1}\leq\beta$ set $X_{i}=\left(\epsilon V_{i}^{1/p}\right)^{1-U^{\prime}_{1}/\beta}$, otherwise set
          $X_{i}=\left[1-\log(1-U^{\prime}_{1})+\log(1-\beta)\right]^{1/p}.$
        3. (b.3)
          If $U^{\prime}_{2}\leq\phi_{2}(X_{i})$ set $R_{i}=X_{i}$, otherwise go back to (b.1).
    4. 4.
      Set $W_{i}=R_{i}V_{i}^{-1/p}S_{i}$.
  4. IV.
    For any $t\in[0,T]$, set $\tilde{X}^{\epsilon}_{t}=\sum_{i=1}^{N}W_{i}1_{[0,t/T]}(U)$.
We now give a small simulation study to illustrate the performance of this algorithm and the approximation in (14). For simplicity, we focus of the bivariate case with $p=1$, drift $\gamma=0$, $Q_{s}=\delta_{1}$ for each $s\in\mathbb{S}^{d-1}$, and we let $\sigma$ be a discrete uniform probability measure on $n$ evenly spaced points in $\mathbb{S}^{1}$. Specifically, we take $\sigma=\frac{1}{n}\sum_{i=1}^{n}\delta_{s_{i}}$, where $s_{i}=(\cos\theta_{i},\sin\theta_{i})$ with $\theta_{i}=\frac{2\pi}{n}(i-1)$, $i=1,2,\dots,n$. In this case $\lambda^{\epsilon}=k_{\epsilon}(s)=\ell(\epsilon)=\Gamma(0,\epsilon)$ and the Lévy measure in (19) simplifies to
$\nu(B)=\frac{1}{n}\sum_{i=1}^{n}\int_{0}^{\infty}1_{B}(rs_{i})e^{-r}r^{-1}\mathrm{d}r,\ \ \ B\in\mathfrak{B}(\mathbb{R}^{d}).$
Here, $G_{V}(\mathrm{d}v;s)=\delta_{1}(\mathrm{d}v)$, which means that $V\sim G_{V}(\cdot;s)$ if and only if $V=1$ with probability $1$. Next, note that $\sigma_{p}=\sigma$, which means that $\sigma_{p}$ is discrete uniform and we can simulate from it using a standard approach. For concreteness we take $n=30$ to be the number of points in the support of $\sigma_{p}$. To simulate from $G_{R}$ we take the tuning parameter $\beta=1/2$.
Since we are in the bivariate case, we write $X_{t}=(X_{1,t},X_{2,t})$ and $\tilde{X}^{\epsilon}_{t}=(\tilde{X}^{\epsilon}_{1,t},\tilde{X}^{\epsilon}_{2,t})$ to denote the Lévy process and the compound Poisson process of large jumps, respectively, at time $t$. It is readily checked that
$\begin{array}[]{ll}\mathrm{E}\left[\tilde{X}_{1,t}^{\epsilon}\right]=te^{-\epsilon}\frac{1}{n}\sum_{i=1}^{n}\cos\theta_{i},&\mathrm{E}\left[\tilde{X}_{2,t}^{\epsilon}\right]=te^{-\epsilon}\frac{1}{n}\sum_{i=1}^{n}\sin\theta_{i},\\ \mathrm{Var}\left(\tilde{X}_{1,t}^{\epsilon}\right)=t(\epsilon+1)e^{-\epsilon}\frac{1}{n}\sum_{i=1}^{n}\cos^{2}\theta_{i},&\mathrm{Var}\left(\tilde{X}_{2,t}^{\epsilon}\right)=t(\epsilon+1)e^{-\epsilon}\frac{1}{n}\sum_{i=1}^{n}\sin^{2}\theta_{i},\end{array}$
and
$\mathrm{Cov}\left(\tilde{X}_{1,t}^{\epsilon},\tilde{X}_{2,t}^{\epsilon}\right)=t(\epsilon+1)e^{-\epsilon}\frac{1}{n}\sum_{i=1}^{n}\cos\theta_{i}\sin\theta_{i}.$
We can similarly calculate the means, variances, and the covariance on the components of $X_{t}$. The formulas are the same, but with $\epsilon=0$. Note that these quantities scale linearly in $t$.
Refer to caption Refer to caption
(a) (b)
Figure 1: (a) gives a simulated path of $\tilde{X}^{\epsilon}$, the compound Poisson process of large jumps. (b) gives a simulated path of $Y^{1}$, the Multivariate Dickman Lévy process used to approximate the small jumps. (c) gives the simulated path of $X=\tilde{X}^{\epsilon}+\epsilon Y^{1}$. In all plots $\epsilon=0.1$.
In Figure 1 we plot a sample path of the process. First, in Figure 1(a) we plot a path of $\tilde{X}^{\epsilon}$, the compound Poisson process of large jumps, which was simulated using Algorithm 3. Then, in Figure 1(b) we plot a path of $Y^{1}$, the Multivariate Dickman Lévy process used to approximate the small jumps. This was simulated using (16), where we truncate the infinite sum at $10,000$. Finally, in Figure 1(c) we plot a path of $X=\tilde{X}^{\epsilon}+\epsilon Y^{1}$. This path is based on the paths presented in Figures 1(a) and 1(b). In these simulations we take $\epsilon=0.1$.
Next, we performed a small simulation study to understand the error in our approximation. Toward this end we simulated $N=500,000$ paths of the Lévy process $X$. For each time $t$, let $m_{1}(t)$ and $s^{2}_{1}(t)$ and $m_{2}(t)$ and $s^{2}_{2}(t)$ be the sample means and sample variances for the first and second components, respectively, and let $s_{1,2}(t)$ be the sample covariance. Now, set
$\text{ErrMean}_{i}(t)$ $=$ $\frac{\left|\mathrm{E}\left[X_{i,t}\right]-m_{i}(t)\right|}{t},\ \ i=1,2$
$\text{ErrVar}_{i}(t)$ $=$ $\frac{\left|\mathrm{Var}\left(X_{i,t}\right)-s^{2}_{i}(t)\right|}{t},\ \ i=1,2$
$\text{ErrCov}(t)$ $=$ $\frac{\left|\mathrm{Cov}\left(X_{1,t},X_{2,t}\right)-s_{1,2}(t)\right|}{t}$
to be the errors in our estimates. Note that we divide these by $t$ since the theoretical values scale linearly in $t$. We then combine these into one total error term given by
$\mathrm{TotalError}(t)=\left(\mathrm{ErrMean}_{1}(t)^{2}+\mathrm{ErrMean}_{2}(t)^{2}+\mathrm{ErrVar}_{1}(t)^{2}+\mathrm{ErrVar}_{2}(t)^{2}+\mathrm{ErrCov}(t)^{2}\right)^{1/2}.$
Further, to understand the performance of Algorithm 3 in simulating $\tilde{X}^{\epsilon}$, we performed a similar simulation study. We again used $N=500,000$ paths and quantified the error analogously, but now using the formulas for the means, variances, and the covariance that are appropriate for this process.
The results of these simulations are presented in Figure 2. In Figure 2(a), we can see that the error in simulating $\tilde{X}^{\epsilon}$ is small for all $\epsilon$’s considered. This not not surprising as Algorithm 3 is exact for all choices of $\epsilon$. In Figure 2(b) we see that our approximate method for simulating the process $X$ works well for small $\epsilon$. In Figure 2(c) we fix $\epsilon=0.1$ and compare the performance of our approach of taking $X\approx\tilde{X}^{\epsilon}+\epsilon Y^{1}$ against a potential approach of just removing the small jumps and taking $X\approx\tilde{X}^{\epsilon}$. We can see that the approach where we model the small jumps using a Dickman Lévy process has significantly less error.
Refer to caption Refer to caption
(a) (b)
Figure 2: Plots of errors. In (a) we give the error when simulating $\tilde{X}^{\epsilon}$, the process of large jumps. In (b) we give the error when simulating $X$, the Lévy process of interest. In (c) we compare the error when simulating $X$ by just the process of large jumps with the sum of the process of large jumps and the Dickman approximation to the small jumps. Here we take $\epsilon=0.1$. All plots are based on $N=500,000$ Monte-Carlo replications.
6 Extension to the BNS Stochastic Volatility Model
In which our approximation results are extended to the Barndorff-Nielsen and Shephard (BNS) stochastic volatility model.
The BNS model is one of the best known continuous time stochastic volatility models and is often used to model financial returns. It was introduced by Barndorff-Nielsen and Shephard in [5]. Our discussion mainly follows Chapter 15 in [9]. The results in this section do not aim for generality, but to demonstrate an approach to deriving approximation results that can be applied in a variety of other contexts. In this section, unless otherwise specified, all processes are assumed to be one dimensional.
Let $\{S_{t}:t\geq 0\}$ be the stock price process. It is typically assumed that $S_{t}=S_{0}e^{X_{t}}$, where $S_{0}$ is the (known) price at time 00 and $\{X_{t}:t\geq 0\}$ is the process of log returns. The BNS model assumes that
$\mathrm{d}X_{t}$ $=$ $\beta h_{t}\mathrm{d}t+\sqrt{h_{t}}\mathrm{d}W_{t}+\rho\mathrm{d}Z_{t},\ \ \ X_{0}=0$
$\mathrm{d}h_{t}$ $=$ $-ch_{t}\mathrm{d}t+\mathrm{d}Z_{t},\ \ \ h_{0}\geq 0,$
where $\beta>0$, $\rho\in\mathbb{R}$, $c>0$, $W=\{W_{t}:t\geq 0\}$ is standard Brownian motion, and $Z=\{Z_{t}:t\geq 0\}$ is a Lévy process independent of $W$ with $Z_{1}\sim\mathrm{ID}_{0}(\nu,0)$ and $\nu((-\infty,0])=0$. We refer to $Z$ as the background driving Lévy process (BDLP). The process $\{h_{t}:t\geq 0\}$ is called the volatility process and our assumptions ensure that $h_{t}\geq 0$ for each $t\geq 0$. We typically take $\rho\leq 0$, which models the so-called leverage effect. This is the empirically observed fact that prices tend to fall when volatility increases, see [5], [9], and the references therein. The system of stochastic differential equations implies that
$X_{t}$ $=$ $\beta\int_{0}^{t}h_{s}\mathrm{d}s+\int_{0}^{t}\sqrt{h_{t}}\mathrm{d}W_{s}+\rho Z_{t}$
and
$h_{t}$ $=$ $h_{0}-c\int_{0}^{t}h_{s}\mathrm{d}s+Z_{t}.$
The process $\{h_{t}:t\geq 0\}$ is an OU-process, see, e.g., [9] or [24]. It can be equivalently represented as
$h_{t}$ $=$ $h_{0}e^{-ct}+\int_{0}^{t}e^{-c(t-s)}\mathrm{d}Z_{s}.$
Equating the two representations of $h_{t}$ and solving gives the useful identity
$\int_{0}^{t}h_{s}\mathrm{d}s$ $=$ $h_{0}\frac{1-e^{-ct}}{c}+\int_{0}^{t}\frac{1-e^{-c(t-s)}}{c}\mathrm{d}Z_{s}.$ (21)
It follows that
$X_{t}$ $=$ $\beta h_{0}\frac{1-e^{-ct}}{c}+\int_{0}^{t}\left(\beta\frac{1-e^{-c(t-s)}}{c}+\rho\right)\mathrm{d}Z_{s}+\int_{0}^{t}\sqrt{h_{s}}\mathrm{d}W_{s}.$
Now consider the truncated version of the process $Z$ given by $Z^{\epsilon}=\{Z^{\epsilon}_{t}:t\geq 0\}$ with $Z_{1}^{\epsilon}\sim\mathrm{ID}_{0}(\nu^{\epsilon},0)$ and the truncated BNS model given by
$\mathrm{d}X_{t}^{\epsilon}$ $=$ $\beta h_{t}^{\epsilon}\mathrm{d}t+\sqrt{h_{t}^{\epsilon}}\mathrm{d}W_{t}+\rho\mathrm{d}Z_{t}^{\epsilon}\ \ \ X^{\epsilon}_{0}=0$
$\mathrm{d}h_{t}^{\epsilon}$ $=$ $-ch_{t}^{\epsilon}\mathrm{d}t+\mathrm{d}Z_{t}^{\epsilon},\ \ \ h_{0}^{\epsilon}=\epsilon h_{0},$
where $Z^{\epsilon}$ is independent of $W$. We will show that for small $\epsilon>0$ this can be approximated by a BNS model with a Dickman BDLP. Specifically, let $Z^{*}=\{Z^{*}_{t}:t\geq 0\}$ be a Lévy process with $Z^{*}_{1}\sim\mathrm{MD}^{1}(\sigma)$. Assume that $Z^{*}$ is independent of $W$ and take
$\mathrm{d}X_{t}^{*}$ $=$ $\beta h_{t}^{*}\mathrm{d}t+\sqrt{h_{t}^{*}}\mathrm{d}W_{t}+\rho\mathrm{d}Z_{t}^{*}\ \ \ X_{0}^{*}=0$
$\mathrm{d}h_{t}^{*}$ $=$ $-ch_{t}^{*}\mathrm{d}t+\mathrm{d}Z_{t}^{*},\ \ \ h_{0}^{*}=h_{0}.$
We now give the joint convergence for the various components of the model. We note that here different components have different rates of convergence.
Proposition 3.
Fix $t>0$ and let $V_{t}^{\epsilon}=(\frac{1}{\epsilon}\int_{0}^{t}h_{s}^{\epsilon}\mathrm{d}s,\frac{1}{\sqrt{\epsilon}}\int_{0}^{t}\sqrt{h_{s}^{\epsilon}}\mathrm{d}W_{s},\frac{1}{\epsilon}Z_{t}^{\epsilon})$ and $V^{*}_{t}=(\int_{0}^{t}h_{s}^{*}\mathrm{d}s,$ $\int_{0}^{t}\sqrt{h_{s}^{*}}\mathrm{d}W_{s},Z_{t}^{*})$. If any of the equivalent conditions in Proposition 1 hold, then for any $t\geq 0$, we have $V_{t}^{\epsilon}\stackrel{{\scriptstyle d}}{{\rightarrow}}V_{t}^{*}$ as $\epsilon\downarrow 0$.
We note that this only gives convergence at one $t$ at a time. Extending it to a stronger form of convergence is an important direction for future work. The result suggests that, when the assumptions hold, for small $\epsilon>0$, we have
$X_{t}^{\epsilon}$ $\stackrel{{\scriptstyle d}}{{\approx}}$ $\epsilon\beta\int_{0}^{t}h_{s}^{*}\mathrm{d}s+\sqrt{\epsilon}\int_{0}^{t}\sqrt{h_{s}^{*}}\mathrm{d}W_{s}+\epsilon\rho Z_{t}^{*}.$
7 Proofs
In which the proofs are given.
7.1 Proofs for Section 2
Lemma 2.
Let $Z=\{Z_{t}:t\geq 0\}$ be a Lévy process on $\mathbb{R}^{d}$, let $a\in\mathbb{R}$, let $z_{1},z_{2}\in\mathbb{R}^{\ell}$, and let $F$, $F^{(1)},F^{(2)}$ be integrable kernels on ×$\mathbb{R}^{\ell\times d}$. For each $t\geq 0$, we have
$\int_{0}^{\infty}\left(z_{1}^{\top}F^{(1)}(s,t)+z_{2}^{\top}F^{(2)}(s,t)\right)\mathrm{d}Z_{s}=z_{1}^{\top}\int_{0}^{\infty}F^{(1)}(s,t)\mathrm{d}Z_{s}+z_{2}^{\top}\int_{0}^{\infty}F^{(2)}(s,t)\mathrm{d}Z_{s}\ \ a.s.$
and
$a\int_{0}^{\infty}F(s,t)\mathrm{d}Z_{s}=\int_{0}^{\infty}F(s,t)\mathrm{d}Z^{\prime}_{s}\ \ a.s.,$
where $Z^{\prime}=\{Z^{\prime}_{t}:t\geq 0\}$ is a Lévy process with $Z^{\prime}_{t}=aZ_{t}$ for every $t\geq 0$.
Proof.
The fact that $Z^{\prime}$ is a Lévy process follows from Proposition 11.10 in [27]. The rest follows immediately from Definitions 2.14, 2.16, and 2.20 and Proposition 2.15 in [28]. ∎
Proof of Theorem 1.
In this proof all limits are as $n\to\infty$. In light of the second part of Lemma 2 and without loss of generality, we assume that $c_{n}=1$ for each $n$. Fix $h^{\prime}>\left(1\vee h\right)$ such that $M^{(0)}(\{x:|x|=h^{\prime}\})=0$. Let $b_{h^{\prime}}^{(n)}=b^{(n)}+\int_{1<|x|\leq h^{\prime}}xM^{(n)}(\mathrm{d}x)$ and $A_{h^{\prime}}^{(n)}=A^{(n)}+\int_{|x|\leq h^{\prime}}xx^{\top}M^{(n)}(\mathrm{d}x)$. Since $Z_{1}^{(n)}\stackrel{{\scriptstyle d}}{{\rightarrow}}Z_{1}^{(0)}$, Theorem 15.14 in [16] implies that $M^{(n)}\stackrel{{\scriptstyle v}}{{\rightarrow}}M^{(0)}$, $A^{(n)}_{h^{\prime}}\to A^{(0)}_{h^{\prime}}$, and $b_{h^{\prime}}^{(n)}\to b_{h^{\prime}}^{(0)}$.
Fix $N\in\mathbb{N}$, let $0\leq t_{1}<t_{2}<\dots<t_{N}<\infty$, let $z_{1},z_{2},\dots,z_{N}\in\mathbb{R}^{\ell}$, and set
$Y^{(n)}=z_{1}^{\top}X^{(n)}_{t_{1}}+z_{2}^{\top}X^{(n)}_{t_{2}}+\dots+z_{N}^{\top}X^{(n)}_{t_{N}}.$
Note that $Y^{(n)}$ is an $\mathbb{R}$-valued random variable. It suffices to show that $\mathrm{E}[e^{iY^{(n)}}]\to\mathrm{E}[e^{iY^{(0)}}]$. By Lemma 2, we have
$Y^{(n)}=\int_{0}^{\infty}F_{*}(s)\mathrm{d}Z^{(n)}_{s}\in\mathbb{R},$
where
$F_{*}(s)=z_{1}^{\top}F(s,t_{1})+z_{1}^{\top}F(s,t_{2})+\dots+z_{N}^{\top}F(s,t_{N})\in\mathbb{R}^{1\times d}.$
The cgf of $Y^{(n)}$ is given by
$\int_{0}^{\infty}C_{\mu^{(n)}}(F^{\top}_{*}(s)z)\mathrm{d}s,\ \ \ z\in\mathbb{R},$ (22)
where $C_{\mu^{(n)}}$ is the cgf of $\mu^{(n)}$ and is of the form given in (1). Since $Z_{1}^{(n)}\stackrel{{\scriptstyle d}}{{\rightarrow}}Z_{1}^{(0)}$, we have $C_{\mu^{(n)}}(F^{\top}_{*}(s)z)\to C_{\mu^{(0)}}(F^{\top}_{*}(s)z)$ for each $z\in\mathbb{R}$ and $s\geq 0$. We need to show convergence of the cgf of $Y^{(n)}$ at $z=1$. It suffices to show that we can use dominated convergence.
For $z=1$, any $s\geq 0$, and taking into account the definition of $b_{h^{\prime}}$, we have
$\left|C_{\mu^{(n)}}(F^{\top}_{*}(s))\right|$
$=\left|-\langle F^{\top}_{*}(s),A^{(n)}F^{\top}_{*}(s)\rangle+i\left\langle b^{(n)}_{h^{\prime}},F^{\top}_{*}(s)\right\rangle\right.$
$\qquad\qquad\qquad\qquad\qquad\qquad\qquad\left.+\int_{\mathbb{R}^{d}}\left(e^{i\left\langle F^{\top}_{*}(s),x\right\rangle}-1-i\left\langle F^{\top}_{*}(s),x\right\rangle 1_{[|x|\leq h^{\prime}]}\right)M^{(n)}(\mathrm{d}x)\right|$
$\leq\langle F^{\top}_{*}(s),A^{(n)}F^{\top}_{*}(s)\rangle+\left|b^{(n)}_{h^{\prime}}\right|\left|F^{\top}_{*}(s)\right|+\int_{|x|\leq h^{\prime}}\left\langle F^{\top}_{*}(s),x\right\rangle^{2}M^{(n)}(\mathrm{d}x)$
$\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad+\int_{|x|>h^{\prime}}\left(\left|\langle F^{\top}_{*}(s),x\rangle\right|\wedge 2\right)M^{(n)}(\mathrm{d}x)$
$=\langle F^{\top}_{*}(s),A_{h^{\prime}}^{(n)}F^{\top}_{*}(s)\rangle+\left|b^{(n)}_{h^{\prime}}\right|\left|F^{\top}_{*}(s)\right|+\int_{|x|>h^{\prime}}\left(\left|\langle F^{\top}_{*}(s),x\rangle\right|\wedge 2\right)M^{(n)}(\mathrm{d}x)$
$\leq\|A_{h^{\prime}}^{(n)}\||F^{\top}_{*}(s)|^{2}+\left|b^{(n)}_{h^{\prime}}\right|\left|F^{\top}_{*}(s)\right|+\min\left\{|F^{\top}_{*}(s)|\int_{|x|>h^{\prime}}|x|M^{(n)}(\mathrm{d}x),2M^{(n)}(\{x:|x|>h^{\prime}\})\right\}$
$\leq\|A_{h^{\prime}}^{(n)}\||F^{\top}_{*}(s)|^{2}+\left|b^{(n)}_{h^{\prime}}\right|\left|F^{\top}_{*}(s)\right|+\min\left\{|F^{\top}_{*}(s)|\int_{|x|>h}|x|M^{(n)}(\mathrm{d}x),2M^{(n)}(\{x:|x|>h^{\prime}\})\right\},$
[…]
$\langle F^{\top}_{*}(s),A^{(n)}_{h^{\prime}}F^{\top}_{*}(s)\rangle=\langle F^{\top}_{*}(s),A^{(n)}F^{\top}_{*}(s)\rangle+\int_{|x|\leq h^{\prime}}\langle F^{\top}_{*}(s),x\rangle^{2}M^{(n)}(\mathrm{d}x),$
the bounds $|e^{ix}-1|\leq|x|\wedge 2$ and $|e^{ix}-1-x|\leq x^{2}$ for any $x\in\mathbb{R}$, see e.g. (26.4) in [7], and various facts about inner products and matrix norms that can be found in, e.g., Section 2.1 of [19]. It follows that, under (i), we have
$\left|C_{\mu^{(n)}}(F^{\top}_{*}(s))\right|$$\leq$ $\left(|F^{\top}_{*}(s)|^{2}\vee|F^{\top}_{*}(s)|\right)\left(\left\|A_{h^{\prime}}^{(n)}\right\|+\left|b^{(n)}_{h^{\prime}}\right|+\int_{|x|>h}|x|M^{(n)}(\mathrm{d}x)\right).$
The limsup of the second term is $\|A_{h^{\prime}}^{(0)}\|+|b^{(0)}_{h^{\prime}}|+M_{h},$ and the result follows by (4). The case when (ii) holds is similar so we only sketch the proof. The main difference is that, in this case, the integral in (22) is over a bounded set and that we upper bound the minimum by $2M^{(n)}(\{x:|x|>h^{\prime}\})$. We then use the fact that, by the Portmanteau Theorem in [2], $M^{(n)}\stackrel{{\scriptstyle v}}{{\rightarrow}}M^{(0)}$ implies that $M^{(n)}(\{x:|x|>h^{\prime}\})\to M^{(0)}(\{x:|x|>h^{\prime}\})$. ∎
7.2 Proofs for Section 3
The proof of Theorem 2 follows from two lemmas, which may be of independent interest.
Lemma 3.
Let $D$ be Borel measure on $\mathbb{R}^{d}$ such that $D(\{0\})=0$ and for every $0<m<M<\infty$
$\int_{m<|x|<M}D(\mathrm{d}x)<\infty.$
If, for every $a>0$ and every $B\in\mathfrak{B}(\mathbb{R}^{d})$ that is bounded away from 00 and infinity, we have
$D(aB)=D(B),$ (23)
then there exists a finite Borel measure $\sigma$ on $\mathbb{S}^{d-1}$ with
$D(B)=\int_{\mathbb{S}^{d-1}}\int_{0}^{\infty}1_{B}(rs)r^{-1}\mathrm{d}r\sigma(\mathrm{d}s),\qquad B\in\mathfrak{B}(\mathbb{R}^{d}).$ (24)
Proof.
Let $\sigma$ be the finite Borel measure on $\mathbb{S}^{d-1}$ given by
$\sigma(C)=D((1,e]C),\ \ C\in\mathfrak{B}(\mathbb{S}^{d-1}).$
Fix $C\in\mathfrak{B}(\mathbb{S}^{d-1})$. For any positive integers $m$ and $n$, (23) implies that
$\frac{n}{m}\sigma(C)$ $=$ $\frac{n}{m}\sum_{i=1}^{m}D((e^{(i-1)/m},e^{i/m}]C)=\frac{n}{m}\sum_{i=1}^{m}D(e^{(i-1)/m}(1,e^{1/m}]C)$
$=$ $nD((1,e^{1/m}]C)=\sum_{i=1}^{n}D((e^{(i-1)/m},e^{i/m}]C)=D((1,e^{n/m}]C).$
Applying (23) again gives
$D((e^{-n/m},1]C)=D((1,e^{n/m}]C)=\frac{n}{m}\sigma(C).$
By the continuity of measures (see Theorem 10.2 in [7]) and the fact that rational numbers are dense in $\mathbb{R}$, it follows that for $y>0$ we have
$D((1,e^{y}]C)=D((e^{-y},1]C)=\sigma(C)y.$
Equivalently for $x>1$, we have
$D((1,x]C)=D((1/x,1]C)=\sigma(C)\log x.$
Noting that for $x>1$
$\log x=\int_{1}^{x}r^{-1}\mathrm{d}r=\int_{1/x}^{1}r^{-1}\mathrm{d}r$
allows us to conclude that, for any $0<a<b<\infty$,
$D((a,b]C)=\int_{\mathbb{S}^{d-1}}\int_{0}^{\infty}1_{(a,b]C}(rs)r^{-1}\mathrm{d}r\sigma(\mathrm{d}s).$
Now let $\mathscr{A}=\{(a,b]C:C\in\mathfrak{B}(\mathbb{S}^{d-1}),0<a<b\}\cup\{\{0\},\emptyset\}$ and note that this is a $\pi$-system satisfying $\sigma(\mathscr{A})=\mathfrak{B}(\mathbb{R}^{d})$. From here the result follows by Theorem 10.3 in [7]. ∎
Lemma 4.
Let $L$ be a non-zero Borel measure on $\mathbb{R}^{d}$ such that $L^{\epsilon}(\mathrm{d}x)=1_{[|x|\leq\epsilon]}L(\mathrm{d}x)$ is a Lévy measure for each $\epsilon>0$. Assume that there exists a function $\psi:(0,\infty)\mapsto[0,\infty)$ such that for any $\epsilon,\gamma>0$
$L^{\epsilon}(\psi(\gamma)B)=L^{\epsilon/\gamma}(B),\qquad B\in\mathfrak{B}(\mathbb{R}^{d}).$ (25)
Then $\psi(\gamma)=\gamma$ and there exists a finite Borel measure $\sigma$ on $\mathbb{S}^{d-1}$ such that $L$ is of the form given in (24).
Proof.
We begin by verifying the following properties of $\psi$ and $L$:
(a) $\psi(\gamma)>0$ for every $\gamma>0$ and $L^{\epsilon}(\mathbb{R}^{d})>0$ for every $\epsilon>0$;
(b) if there are $\epsilon,\gamma,\beta_{1},\beta_{2}>0$ with $L^{\epsilon}(\beta_{1}B)=L^{\epsilon}(\beta_{2}B)=L^{\epsilon/\gamma}(B)$ for every $B\in\mathfrak{B}(\mathbb{R}^{d})$, then $\beta_{1}=\beta_{2}$;
(c) for $\gamma,\gamma^{\prime}>0$, we have $\psi(\gamma\gamma^{\prime})=\psi(\gamma)\psi(\gamma^{\prime})$, $\psi(1)=1$, and, $\psi(1/\gamma)=\frac{1}{\psi(\gamma)}$;
(d) $\psi$ is a continuous function.
First note that, in the presence of (a) and (b), property (c) follows immediately from (25). Next, we turn to (a). If $\psi(\gamma)=0$ for some $\gamma>0$, then for any $\epsilon>0$ we have $L^{\epsilon}(\mathbb{R}^{d})=L^{\epsilon\gamma}(\psi(\gamma)\mathbb{R}^{d})=L^{\epsilon\gamma}(\{0\})=0$, which implies that $L=0$. Similarly, if $L^{\epsilon}(\mathbb{R}^{d})=0$ for some $\epsilon>0$, then for any $\epsilon^{\prime}>0$ we have $\psi(\epsilon/\epsilon^{\prime})\mathbb{R}^{d}=\mathbb{R}^{d}$ and thus $0=L^{\epsilon}(\mathbb{R}^{d})=L^{\epsilon}(\psi(\epsilon/\epsilon^{\prime})\mathbb{R}^{d})=L^{\epsilon^{\prime}}(\mathbb{R}^{d})$, so $L^{\epsilon^{\prime}}=0$ and hence $L=0$.
We now turn to (b). Assume for the sake of contradiction that there are $\gamma,\epsilon>0$ and constants $0<\beta_{1}<\beta_{2}<\infty$, with $L^{\epsilon}(\beta_{1}B)=L^{\epsilon}(\beta_{2}B)=L^{\epsilon/\gamma}(B)$ for each $B\in\mathfrak{B}(\mathbb{R}^{d})$. Let $\rho=\beta_{2}/\beta_{1}$ and note that $\rho>1$. Next, note that $L^{\epsilon}(B)=L^{\epsilon}(\frac{\beta_{1}}{\beta_{1}}B)=L^{\epsilon/\gamma}(\frac{1}{\beta_{1}}B)$ and thus that $L^{\epsilon}(\rho B)=L^{\epsilon/\gamma}(\frac{1}{\beta_{1}}B)=L^{\epsilon}(B)$. It follows that $L^{\epsilon}(\rho^{n}B)=L^{\epsilon}(B)$ for each $n\in\mathbb{N}$. Now assume that $B$ is bounded away from 00 and satisfies $L^{\epsilon}(B)>0$; such a $B$ exists by (a). It follows that $\rho^{n}B\to\emptyset$ and, by the continuity of measures (see e.g. Problem 10.4 in [7]) and the fact that $L^{\epsilon}$ is finite away from 00,
$0<L^{\epsilon}(B)=\lim_{n\to\infty}L^{\epsilon}(\rho^{n}B)=0,$
which is a contradiction.
Turning to (d), let $\gamma_{n}\to\gamma\in(0,\infty)$ and let $B\in\mathfrak{B}(\mathbb{R}^{d})$ be any set that is bounded away from 00 and satisfies $L^{1/\gamma}(B)>0$. By dominated convergence
$\lim_{n\to\infty}L^{1}(\psi(\gamma_{n})B)=\lim_{n\to\infty}L^{1/\gamma_{n}}(B)=\lim_{n\to\infty}\int_{|x|\leq 1/\gamma_{n}}1_{B}(x)L(\mathrm{d}x)=L^{1/\gamma}(B)>0.$
Now, assume that $\psi(\gamma_{n})\to\beta\in[1,\infty]$ (possibly along a subsequence). Here, we take $\beta\geq 1$ without loss of generality by (c). Then (along the same subsequence) $\lim_{n\to\infty}\psi(\gamma_{n})B=\beta B$ if $\beta<\infty$ and otherwise the limit is $\emptyset$. By the continuity of measures, it follows that if $\beta<\infty$
$0<L^{1/\gamma}(B)=\lim_{n\to\infty}L^{1}(\psi(\gamma_{n})B)=L^{1}(\beta B)$
and the limit equals 00 if $\beta=\infty$. Since it cannot be 00, we have $\beta<\infty$ and, by (b), we have $\beta=\psi(\gamma)$, which gives (d).
Now let $\phi(x)=\psi(e^{x})$ for $x\in\mathbb{R}$. By (d) $\phi$ is continuous and by (c) it satisfies $\phi(x+y)=\phi(x)\phi(y)$ for $x,y\in\mathbb{R}$. It follows, see e.g. page 197 in [26], that $\phi(x)=e^{cx}$ for some $c\in\mathbb{R}$. Thus, $\psi(\gamma)=\gamma^{c}$. We now show that $c>0$. If $c=0$, then $L^{\epsilon_{1}}(B)=L^{\epsilon_{2}}(B)$ for every $\epsilon_{1},\epsilon_{2}>0$, which means that $L=0$. If $c<0$, then for $\gamma\in(0,1)$ we have $\psi(\gamma^{n})=\gamma^{nc}\to\infty$. Let $B\in\mathfrak{B}(\mathbb{R}^{d})$ be any set with that is bounded away from 00 and satisfies $L(B)>0$. We have $\psi(\gamma^{n})B\to\emptyset$ and thus
$\lim_{n\to\infty}L^{1}(\psi(\gamma^{n})B)=\lim_{n\to\infty}L^{1}(\emptyset)=0.$
On the other hand, by monotone convergence
$\lim_{n\to\infty}L^{1}(\psi(\gamma^{n})B)=\lim_{n\to\infty}L^{1/\gamma^{n}}(B)=\lim_{n\to\infty}\int_{|x|\leq 1/\gamma^{n}}1_{B}(x)L(\mathrm{d}x)=L(B)>0,$
which gives the contradiction.
We have now shown that there exists a $c>0$ such that for every $\gamma>0$ we have $\psi(\gamma)=\gamma^{c}$. From here the results follows from Lemma 3. We just need to show that $L$ satisfies (23). Fix $a>0$ and let $B$ be a Borel set that is bounded away from infinity. There exists an $\epsilon>0$ such that $|x|<\min\{\epsilon,a^{1-1/c}\epsilon\}$ for every $x\in B$. Thus, $L(B)=L^{\epsilon}(B)$. It follows that
$L(aB)=L^{a\epsilon}(aB)=L^{a^{1-1/c}\epsilon}(B)=L(B).$
The fact that $c=1$ follows from the fact that this is the only $c$ that works for the measure in (24). ∎
Proof of Theorem 2.
It is readily checked that for any $\epsilon,\gamma>0$ we have $\frac{1}{\psi(\gamma)}X^{\epsilon}\sim\mathrm{ID}(0,M^{\epsilon}_{*},b_{*}(\epsilon,\gamma))$, with
$M_{*}^{\epsilon}(B)=\int_{|x|\leq\epsilon}1_{B}\left(\frac{1}{\psi(\gamma)}x\right)M(\mathrm{d}x),\quad B\in\mathfrak{B}(\mathbb{R}^{d})$
and
$b_{*}(\epsilon,\gamma)=\frac{1}{\psi(\gamma)}\int_{|x|\leq\epsilon}x\left(1_{[{|x|\leq\psi(\gamma)]}}-1_{[|x|\leq 1]}\right)M(\mathrm{d}x).$
Thus, for every $\epsilon,\gamma>0$ and all $B\in\mathfrak{B}(\mathbb{R}^{d})$
$M^{\epsilon/\gamma}(B)=M_{*}^{\epsilon}(B)=M^{\epsilon}(\psi(\gamma)B).$
This satisfies the assumption of Lemma 4. Hence, $\psi(\gamma)=\gamma$ and $M$ is of the required form. ∎
Proof of Theorem 3.
For any $\tau>0$ and any integer $i\geq 1$ we have
$\Gamma_{N(\tau)+i}=\tau+\Gamma_{i}^{\prime},$ (26)
where $\Gamma_{i}^{\prime}=\sum_{k=1}^{i}E_{k}^{\prime}$, $E^{\prime}_{k}=E_{N(\tau)+k}$ for $k\geq 2$ and $E^{\prime}_{1}=\Gamma_{N(\tau)+1}-\tau$. By the memoryless property of the exponential distribution, $E_{1}^{\prime}\sim\mathrm{Exp}(\theta)$. If $a=0$, then, with probability $1$, $X_{at}=X_{0}=0=0*X_{t}$. Next, for any $a\in(0,1]$
$X_{at}=a\sum_{i=1}^{\infty}1_{[\Gamma_{i}-\log(1/a)>\log(1/t)]}e^{-\left(\Gamma_{i}-\log(1/a)\right)}\xi_{i}=a\sum_{i=1}^{\infty}1_{[\Gamma_{i}^{\prime}>\log(1/t)]}e^{-\Gamma_{i}^{\prime}}\xi_{i},$
where we use (26) with $\tau=\log(1/a)$. From here, the fact that $\{\Gamma_{i}^{\prime}:i=1,2,\dots\}\stackrel{{\scriptstyle d}}{{=}}\{\Gamma_{i}:i=1,2,\dots\}$ gives the self-similarity. Now to show that the random variables have the correct distributions. We have $X_{1}\sim\mathrm{MD}(\sigma)$ by (11). Next, from the self-similarity, we have $X_{t}\stackrel{{\scriptstyle d}}{{=}}tX_{1}\sim\mathrm{MD}^{t}(\sigma)$ by (9). ∎
7.3 Proofs for Section 4
Proof of Proposition 1.
The equivalence between Conditions 1 and 2 follows easily from Lemma 4.9 in [12]. We just note that for every $h>0$ we have $D^{1}(\{x\in\mathbb{R}^{d}:|x|=h\})=0$,
$\nu((\epsilon h,\epsilon]C)=M^{\epsilon}\left(\left\{x\in\mathbb{R}^{d}:|x|>h,\frac{x}{|x|}\in C\right\}\right),$
and
$\sigma(C)\log\frac{1}{h}=D^{1}\left(\left\{x\in\mathbb{R}^{d}:|x|>h,\frac{x}{|x|}\in C\right\}\right).$
We now show that Condition 1 implies Condition 3. First fix $p,\epsilon>0$, $N\in\mathbb{N}$, $C\in\mathfrak{B}(\mathbb{S}^{d-1})$ with $\sigma(\partial C)=0$, and let $\epsilon_{k}=2^{-(k-1)}\epsilon$. We have
$\int_{(2^{-N},1]C}|x|^{p}M^{\epsilon}(\mathrm{d}x)$ $=$ $\frac{1}{\epsilon^{p}}\int_{(\frac{\epsilon}{2^{N}},\epsilon]C}|x|^{p}\nu(\mathrm{d}x)$
$=$ $\sum\limits_{k=1}^{N}\frac{1}{\epsilon^{p}}\int_{\big{(}\frac{\epsilon}{2^{k}},\frac{\epsilon}{2^{k-1}}\big{]}C}|x|^{p}\nu(\mathrm{d}x)$
$=$ $\sum\limits_{k=1}^{N}\frac{1}{2^{p(k-1)}\epsilon^{p}_{k}}\int_{\big{(}\frac{1}{2}\epsilon_{k},\epsilon_{k}\big{]}C}|x|^{p}\nu(\mathrm{d}x)$
$=$ $\sum\limits_{k=1}^{N}2^{-p(k-1)}\int_{(\frac{1}{2},1]C}|x|^{p}M^{\epsilon_{k}}(\mathrm{d}x).$
By the version of the Portmanteau Theorem given in [2], for any $h\in(0,1)$ we have
$\int_{(h,1]C}|x|^{p}M^{\epsilon}(\mathrm{d}x)\to\int_{(h,1]C}|x|^{p}D^{1}(\mathrm{d}x)=\int_{C}\int_{h}^{1}r^{p-1}\mathrm{d}r\sigma(\mathrm{d}s)=\frac{\sigma(C)}{p}(1-h^{p})$
as $\epsilon\downarrow 0$. It follows that, for every $\theta>0$ there exists a $\delta>0$ such that if $0<\epsilon<\delta$, then
$\left|\int_{(h,1]C}|x|^{p}M^{\epsilon}(\mathrm{d}x)-\frac{\sigma(C)}{p}(1-h^{p})\right|\leq\frac{\theta}{\sum_{k=1}^{\infty}2^{-p(k-1)}}.$
Taking $h=1/2$, $0<\epsilon<\delta$, noting that $\epsilon_{n}\in(0,\epsilon]$, and applying the triangle inequality gives
$\left|\int_{(2^{-N},1]C}|x|^{p}M^{\epsilon}(\mathrm{d}x)-\sum_{k=1}^{N}\left(\frac{1}{2^{p}}\right)^{k-1}\frac{\sigma(C)}{p}(1-1/2^{p})\right|\leq\theta.$
Now taking the limit as $N\to\infty$ and applying monotone convergence gives
$\left|\frac{1}{\epsilon^{p}}\int_{(0,\epsilon]C}|x|^{p}\nu(\mathrm{d}x)-\frac{\sigma(C)}{p}\right|=\left|\int_{(0,1]C}|x|^{p}M^{\epsilon}(\mathrm{d}x)-\frac{\sigma(C)}{p}\right|\leq\theta.$
Condition 3 is now proved. Note that we did not assume $\frac{1}{\epsilon^{p}}\int_{(0,\epsilon]C}|x|^{p}\nu(\mathrm{d}x)<\infty$. The fact that this is finite is part of the result.
It is immediate that Condition 3 implies Condition 4. We now show that Condition 4 implies Condition 2, which will complete the proof. Define $\eta_{\epsilon}$ and $\eta$ to be finite Borel measures on $\mathbb{R}^{d}$ such that, for $B\in\mathfrak{B}(\mathbb{R}^{d})$, we have
$\eta_{\epsilon}(B)=\int_{B}|x|^{p}M^{\epsilon}(\mathrm{d}x)\ \mbox{ and }\ \eta(B)=\int_{B}|x|^{p}D^{1}(\mathrm{d}x).$
Fix $C\in\mathfrak{B}(\mathbb{S}^{d-1})$ with $\sigma(\partial C)=0$, $h>0$, and $\epsilon>0$. Let $h_{0}=h\wedge 1$ and note that
$\eta_{\epsilon}\left((0,h]C\right)=\int_{(0,h]C}|x|^{p}M^{\epsilon}(\mathrm{d}x)=\int_{(0,\epsilon h]C}\frac{|x|^{p}}{\epsilon^{p}}\nu^{\epsilon}(\mathrm{d}x)=\frac{1}{\epsilon^{p}}\int_{(0,\epsilon h_{0}]C}|x|^{p}\nu(\mathrm{d}x)$
and
$\eta\left((0,h]C\right)=\int_{(0,h]C}|x|^{p}D^{1}(\mathrm{d}x)=\int_{C}\int_{0}^{h_{0}}|rs|^{p}r^{-1}\mathrm{d}r\sigma(\mathrm{d}s)=h_{0}^{p}\frac{\sigma(C)}{p},$
where we use the fact that $|s|=1$ for every $s\in\mathbb{S}^{d-1}$. Since Condition 4 holds
$\lim\limits_{\epsilon\downarrow 0}\eta_{\epsilon}\left((0,h]C\right)=h_{0}^{p}\lim\limits_{\epsilon\downarrow 0}\frac{1}{h_{0}^{p}\epsilon^{p}}\int_{(0,\epsilon h_{0}]C}|x|^{p}\nu(\mathrm{d}x)=h_{0}^{p}\frac{\sigma(C)}{p}=\eta\left((0,h]C\right).$
Hence, if $10<h<$ 1
$\eta_{\epsilon}\left((h,\infty)C\right)=\eta_{\epsilon}\left((h,1]C\right)$ $=$ $\eta_{\epsilon}\left((0,1]C\right)-\eta_{\epsilon}\left((0,h]C\right)$
$\rightarrow$ $\eta\left((0,1]C\right)-\eta\left((0,h]C\right)=\eta\left((h,1]C\right)=\eta\left((h,\infty)C\right)$
and if $h\geq 1$
$\eta_{\epsilon}\left((h,\infty)C\right)=0=\eta\left((h,\infty)C\right).$
From here Lemma 4.9 in [12] implies that $\eta_{\epsilon}\stackrel{{\scriptstyle v}}{{\rightarrow}}\eta$ as $\epsilon\downarrow 0$. Next, noting that $\frac{1}{|x|^{p}}$ is bounded and continuous when away from zero and applying the Portmanteau Theorem of [2] show that, for any $h\in(0,1)$,
$\lim_{\epsilon\downarrow 0}\nu((\epsilon h,\epsilon])=\lim_{\epsilon\downarrow 0}\int_{(h,1]C}\frac{1}{|x|^{p}}\eta_{\epsilon}(\mathrm{d}x)=\int_{(h,1]C}\frac{1}{|x|^{p}}\eta(\mathrm{d}x)=D^{1}\left((h,1]C\right)=\sigma(C)\log\frac{1}{h},$
which gives the result. ∎
Proof of Theorem 4.
We only prove the first part, as the second part follows immediately by combining the first part with Slutsky’s Theorem. Theorem 15.17 in [16] implies that $\epsilon^{-1}{X^{\epsilon}}\stackrel{{\scriptstyle d}}{{\rightarrow}}Y^{1}$ if and only if $\epsilon^{-1}{X_{1}^{\epsilon}}\stackrel{{\scriptstyle d}}{{\rightarrow}}Y^{1}_{1}$. By a version of Theorem 15.14 in [16] (see also Theorem of 8.7 of [27] or Theorem 3.1.16 in [19]) we have $\epsilon^{-1}X^{\epsilon}_{1}\stackrel{{\scriptstyle d}}{{\rightarrow}}Y^{1}_{1}$ if and only if the following three conditions hold:
  • 1.
    $M^{\epsilon}\stackrel{{\scriptstyle v}}{{\rightarrow}}D^{1}$ as $\epsilon\downarrow 0$;
  • 2.
    $\lim\limits_{\delta\downarrow 0}\limsup\limits_{\epsilon\downarrow 0}\int_{|x|\leq\delta}\langle z,x\rangle^{2}M^{\epsilon}(\mathrm{d}x)=0$ for every $z\in\mathbb{R}^{d}$;
  • 3.
    $\lim\limits_{\epsilon\downarrow 0}\int_{|x|\leq 1}xM^{\epsilon}(\mathrm{d}x)=\int_{|x|\leq 1}xD^{1}(\mathrm{d}x)$.
Condition 1 is one of the equivalent conditions in Proposition 1. To complete the proof, we will show that it implies Conditions 2 and 3. To show that it implies Condition 2, note that
00 0 $\leq$ $\lim\limits_{\delta\downarrow 0}\limsup_{\epsilon\downarrow 0}\int_{|x|\leq\delta}\langle z,x\rangle^{2}M^{\epsilon}(\mathrm{d}x)$
$\leq$ $\lim\limits_{\delta\downarrow 0}\limsup_{\epsilon\downarrow 0}\int_{|x|\leq\delta}|z|^{2}|x|^{2}M^{\epsilon}(\mathrm{d}x)$
$\leq$ $\lim\limits_{\delta\downarrow 0}\limsup_{\epsilon\downarrow 0}|z|^{2}\delta\int_{|x|\leq 1}|x|M^{\epsilon}(\mathrm{d}x)$
$=$ $|z|^{2}\lim\limits_{\delta\downarrow 0}\limsup_{\epsilon\downarrow 0}\delta\int_{|x|\leq\epsilon}\frac{|x|}{\epsilon}\nu(\mathrm{d}x)$
$=$ $|z|^{2}\sigma(\mathbb{S}^{d-1})\lim\limits_{\delta\downarrow 0}\delta=0,$
where the last line follows by Proposition 1. Condition 3 can be shown in a manner similar to how we showed that Condition 1 implies Condition 3 in the proof of Proposition 1. There are two main differences. First, we take $\int_{C}s\sigma(\mathrm{d}s)$ instead of $\sigma(C)$. Second, we no longer take the norm of $x$ in the integral and must use the dominated convergence theorem instead of monotone convergence when taking the limit in $N$. We can use dominated convergence here since we have assumed that $\int_{|x|\leq 1}|x|\nu(\mathrm{d}x)<\infty$. ∎
Proof of Corollary 1.
Note that for any $C\in\mathfrak{B}(\mathbb{S}^{d-1})$
$\lim_{\epsilon\downarrow 0}\frac{1}{\epsilon}\int_{(0,\epsilon]C}|x|\nu(\mathrm{d}x)$ $=$ $\lim_{\epsilon\downarrow 0}\frac{1}{\epsilon}\int_{0}^{\epsilon}\int_{C}r\rho(r,s)\sigma(\mathrm{d}s)\mathrm{d}r$
$=$ $\lim_{\epsilon\downarrow 0}\frac{1}{\epsilon}\int_{0}^{\epsilon}\int_{C}\left(r\rho(r,s)-g(s)\right)\sigma(\mathrm{d}s)\mathrm{d}r+\lim_{\epsilon\downarrow 0}\frac{1}{\epsilon}\int_{0}^{\epsilon}\int_{C}g(s)\sigma(\mathrm{d}s)\mathrm{d}r$
$=$ $\int_{C}g(s)\sigma(\mathrm{d}s)=\sigma_{g}(C),$
where the last line follows from the fact that $\left|\int_{C}(r\rho(r,s)-g(s))\sigma(\mathrm{d}s)\right|\leq\int_{\mathbb{S}^{d-1}}|r\rho(r,s)-h(s)|\sigma(\mathrm{d}s)\to 0$ as $r\downarrow 0$. From here the result follows by Proposition 1 and Theorem 4. ∎
Proof of Corollary 2.
[…]
for each $\epsilon>0$, and satisfying
$\sup\limits_{t\in[0,T]}\left|\hat{X}_{t}^{\epsilon}-Y^{1}_{t}\right|\stackrel{{\scriptstyle p}}{{\rightarrow}}0\ \mbox{ as }\ \epsilon\downarrow 0$
for each $T>0$. We can (and will) take $\hat{X}^{\epsilon}$ to be independent of $\tilde{X}^{\epsilon}$, possibly on an enlarged probability space, e.g., where $Y^{1}$ and $\hat{X}_{t}^{\epsilon}$ depend on different coordinates from $\tilde{X}^{\epsilon}$. Next, for $t>0$ set
$Z^{\epsilon}_{t}=\epsilon\left(\hat{X}_{t}^{\epsilon}-Y^{1}_{t}\right)$
and note that (15) holds for each $T>0$. Note further, that $\{Z_{t}:t\geq 0\}$ is a càdlàg process, since it is the difference of two Lévy processes and Lévy processes have càdlàg paths. We have
$X=X^{\epsilon}+\tilde{X}^{\epsilon}+\gamma^{*}\stackrel{{\scriptstyle d}}{{=}}\epsilon\hat{X}^{\epsilon}+\tilde{X}^{\epsilon}+\gamma^{*}=\epsilon Y^{1}+\tilde{X}^{\epsilon}+\gamma^{*}+Z^{\epsilon},$
which completes the proof. ∎
Proof of Theorem 5.
If any of the equivalent condition in Proposition 1 hold, then Theorem 4 implies that $\epsilon^{-1}X^{\epsilon}\stackrel{{\scriptstyle d}}{{\rightarrow}}Y$. From here the result follows by Theorem 1. To use this theorem it suffices to show that assumption (i) holds. To see this, fix $h>1$ and note that for any $\epsilon\in(0,1)$
$\int_{|x|>h}|x|M^{\epsilon}(\mathrm{d}x)=\frac{1}{\epsilon}\int_{|x|>h\epsilon}|x|\nu^{\epsilon}(\mathrm{d}x)=0$
since $\nu^{\epsilon}(\{x:x>h\epsilon\})\leq\nu^{\epsilon}(\{x:x>\epsilon\})=0$. ∎
7.4 Proofs for Section 5
Proof of Proposition 2.
For any $B\in\mathfrak{B}(\mathbb{R}^{d})$, by the tower property of conditional expectation
$\mathrm{P}(W\in B)$ $=$ $\mathrm{E}\left[\mathrm{E}[\mathrm{E}[1_{B}(RV^{-1/p}S)|V,S]|S]\right]$
$=$ $\mathrm{E}\left[\mathrm{E}\left[\frac{1}{\ell(V^{1/p}\epsilon)}\int_{\epsilon V^{1/p}}^{\infty}1_{B}(rV^{-1/p}S)r^{-1}e^{-r^{p}}\mathrm{d}r\Bigm{|}S\right]\right]$
$=$ $\mathrm{E}\left[\mathrm{E}\left[\frac{1}{\ell(V^{1/p}\epsilon)}\int_{\epsilon}^{\infty}1_{B}(rS)r^{-1}e^{-r^{p}V}\mathrm{d}r\Bigm{|}S\right]\right]$
$=$ $\mathrm{E}\left[\int_{0}^{\infty}\frac{1}{\ell(v^{1/p}\epsilon)}\int_{\epsilon}^{\infty}1_{B}(rS)r^{-1}e^{-r^{p}v}\mathrm{d}r\frac{\ell(v^{1/p}\epsilon)}{k_{\epsilon}(S)}Q_{S}(\mathrm{d}v)\right]$
$=$ $\frac{1}{\lambda^{\epsilon}}\int_{\mathbb{S}^{d-1}}\int_{0}^{\infty}\int_{\epsilon}^{\infty}1_{B}(rs)r^{-1}e^{-r^{p}v}\mathrm{d}rQ_{s}(\mathrm{d}v)\sigma(\mathrm{d}s)=\tilde{\nu}^{\epsilon}_{p}(B)$
as required. ∎
Proof of Lemma 1.
It is easily checked that the cumulative distribution function (cdf) corresponding to $h_{1}(\cdot;a,p)$ is $H_{1}(x;a,p)=1-e^{a^{p}-x^{p}}$ for $x>a$. Thus, $H_{1}^{-1}(x;a,p)=\left[a^{p}-\log\left(1-x\right)\right]^{\frac{1}{p}}$ for $x\in[0,1]$. From here, the fact that $U\stackrel{{\scriptstyle d}}{{=}}1-U$ gives the first part.
Next we turn to simulation from $h_{2}$. Note that this distribution is a mixture. With probability $1-\beta$ we must simulate from $h_{1}(\cdot;1,p)$, which we already know how to do. With probability $\beta$ we must simulate from a distribution with pdf $\frac{x^{-1}}{\log(1/a)}1_{[a\leq x<1]}.$ The corresponding cdf is given by $1-\log(x)/\log(a)$ for $a<x<1$. The inverse function is then $a^{1-x}$ for $x\in[0,1]$. To conclude, we note that the conditional distribution of $U/\beta$ given $U\leq\beta$ is $U(0,1)$ and that the conditional distribution of $(1-U)/(1-\beta)$ given $U>\beta$ is $U(0,1)$. ∎
7.5 Proofs for Section 6
Proof of Proposition 3.
To verify the convergence in distribution, we will show convergence of the moment generating functions. We note that, due to the truncation, there is no issue with the existence of these functions. Let $(\mathcal{F}_{t}^{\epsilon})_{t\geq 0}$ be the filtration generated by the Lévy process $\{Z_{t}^{\epsilon}:t\geq 0\}$. For any $u\in\mathbb{R}$, (5) and (21) imply that
$\mathrm{E}\left[e^{u\frac{1}{\sqrt{\epsilon}}\int_{0}^{t}\sqrt{h_{s}^{\epsilon}}\mathrm{d}W_{s}}|\mathcal{F}_{t}^{\epsilon}\right]=e^{\frac{u^{2}}{2\epsilon}\int_{0}^{t}h_{s}^{\epsilon}\mathrm{d}s}=e^{\frac{u^{2}}{2}h_{0}\frac{1-e^{-ct}}{c}}e^{\frac{u^{2}}{2\epsilon}\int_{0}^{t}\frac{1-e^{-c(t-s)}}{c}\mathrm{d}Z^{\epsilon}_{s}}.$
Next, let $U=(u_{1},u_{2},u_{3})$ and note that $Z_{t}^{\epsilon}$ and $\int_{0}^{t}h_{s}^{\epsilon}\mathrm{d}s$ are measurable $\mathcal{F}_{t}$. We have
$\mathrm{E}\left[e^{\langle U,V_{t}^{\epsilon}\rangle}\right]$ $=$ $\mathrm{E}\left[\mathrm{E}\left[e^{u_{1}\frac{1}{\epsilon}\int_{0}^{t}h_{s}^{\epsilon}\mathrm{d}s+u_{2}\frac{1}{\sqrt{\epsilon}}\int_{0}^{t}\sqrt{h_{s}^{\epsilon}}\mathrm{d}W_{s}+u_{3}\frac{1}{\epsilon}Z_{t}^{\epsilon}}|\mathcal{F}_{t}\right]\right]$
$=$ $\mathrm{E}\left[e^{u_{1}\frac{1}{\epsilon}\int_{0}^{t}h_{s}^{\epsilon}\mathrm{d}s+u_{3}\frac{1}{\epsilon}Z_{t}^{\epsilon}}\mathrm{E}\left[e^{u_{2}\frac{1}{\sqrt{\epsilon}}\int_{0}^{t}\sqrt{h_{s}^{\epsilon}}\mathrm{d}W_{s}}|\mathcal{F}_{t}\right]\right]$
$=$ $e^{\left(\frac{u_{2}^{2}}{2}+u_{1}\right)h_{0}\frac{1-e^{-ct}}{c}}\mathrm{E}\left[e^{\frac{1}{\epsilon}\int_{0}^{t}(u_{1}+\frac{u_{2}^{2}}{2})\frac{1-e^{-c(t-s)}}{c}\mathrm{d}Z_{s}^{\epsilon}+\frac{1}{\epsilon}\int_{0}^{t}u_{3}\mathrm{d}Z_{s}^{\epsilon}}\right]$
$=$ $e^{\left(\frac{u_{2}^{2}}{2}+u_{1}\right)h_{0}\frac{1-e^{-ct}}{c}}\mathrm{E}\left[e^{\frac{1}{\epsilon}\int_{0}^{t}\left[(u_{1}+\frac{u_{2}^{2}}{2})\frac{1-e^{-c(t-s)}}{c}+u_{3}\right]\mathrm{d}Z_{s}^{\epsilon}}\right]$
$=$ $e^{\left(\frac{u_{2}^{2}}{2}+u_{1}\right)h_{0}\frac{1-e^{-ct}}{c}}\mathrm{E}\left[e^{\frac{1}{\epsilon}\int_{0}^{t}f_{U}(s)\mathrm{d}Z_{s}^{\epsilon}}\right],$
where $f_{U}(s)=(u_{1}+\frac{u_{2}^{2}}{2})\frac{1-e^{-c(t-s)}}{c}+u_{3}$. By Theorem 5 we have $\frac{1}{\epsilon}\int_{0}^{t}f_{U}(s)\mathrm{d}Z_{s}^{\epsilon}\stackrel{{\scriptstyle d}}{{\rightarrow}}\int_{0}^{t}f_{U}(s)\mathrm{d}Z^{*}_{s}$ as $\epsilon\downarrow 0$. Thus, $\mathrm{E}\left[e^{\langle U,V_{t}^{\epsilon}\rangle}\right]\to e^{\left(\frac{u^{2}}{2}+u_{1}\right)h_{0}\frac{1-e^{-ct}}{c}}\mathrm{E}\left[e^{\int_{0}^{t}f_{U}(s)\mathrm{d}Z^{*}_{s}}\right]$ as $\epsilon\downarrow 0$. To conclude the proof, we note that $e^{\left(\frac{u^{2}}{2}+u_{1}\right)h_{0}\frac{1-e^{-ct}}{c}}\mathrm{E}\left[e^{\int_{0}^{t}f_{U}(s)\mathrm{d}Z^{*}_{s}}\right]=\mathrm{E}\left[e^{\langle U,V^{*}_{t}\rangle}\right]$, which can be shown by arguments similar to those above. ∎